### v1

# Needle: A Generative-AI Powered Monte Carlo Method for Answering Complex Natural Language Queries on Multi-modal Data

Abstract.
Multi-modal data, such as image data sets, often miss the detailed descriptions that properly capture the rich information encoded in them. This makes answering complex natural language queries a major challenge in these domains. In particular, unlike the traditional nearest-neighbor search, where the tuples and the query are modeled as points in a data cube, the query and the tuples are of different natures, making the traditional query answering solutions not directly applicable for such settings. Existing literature addresses this challenge for image data through vector representations jointly trained on natural language and images. This technique, however, underperforms for complex queries due to various reasons.
This paper takes a step towards addressing this challenge by introducing a Generative-AI (GenAI) powered Monte Carlo method that utilizes foundation models to generate synthetic samples that capture the complexity of the natural language query and transform it to the same space of the multi-modal data. Following this method, we develop a system for image data retrieval and propose practical solutions that enable leveraging future advancements in GenAI and vector representations for improving our system’s performance. Our comprehensive experiments on various benchmark datasets verify that our system significantly outperforms state-of-the-art techniques.
Image Databases; Generative AI; Foundation Models;
1. Introduction
a Needle in a haystack: something that is impossible or extremely difficult to find, especially because the area you have to search is too large.”
Cambridge Dictionary
Figure 1. The query results for the queries: (a) [a banana] (row 1), (b) [an unripe banana] (row 2), and (c) [a banana gazing at its reflection in a mirror] (row 3), using Needle (first column) v.s. CLIP (second column).
Multi-modal data, including image datasets, pose new challenges for data management systems. Unlike traditional databases, where tuple attributes are explicitly specified, multi-modal data often miss proper descriptions that capture the rich information encoded in them. As a result, traditional query-answering approaches are not directly applicable in these settings. On the other hand, advanced personal devices such as cell phones have enabled the collection of large volumes of multi-modal data, even at the individual user level. Due to the wealth of information “hidden” in each multi-modal tuple, the users may compose complex natural language queries, describing the tuple(s) they are interested in. To further motivate this, let us consider the following running example.
Example 1.1.
A photo enthusiast has collected a large pool of images over the past several years on her private data repository. She wants to retrieve a specific photo she has in mind. She describes the picture as [A banana gazing at its reflection in a mirror].
Traditional nearest-neighbor (NN-) search algorithms are not directly applicable here since those consider the database tuples and the query as points in a data cube. In contrast, in Example 1.1, (i) each tuple is an image, and (ii) the query is in the form of an unstructured, complex natural language statement.
The state-of-the-art approach for addressing the need for answering natural language queries on image datasets is CLIP (Radford et al., 2021), a vector representation (aka embedding) jointly trained on several hundred million (image, text) pairs. Using CLIP as the embedder, one can transform the query and the dataset images to vector representation in the same space and apply NN-search to answer the query. This approach works satisfactorily for “simple” queries, such as simple object detection tasks. To see a concrete example, following with our running example (Example 1.1), we used a benchmark image dataset (LVIS (Gupta et al., 2019)) and issued the query [A banana]. As reflected in the top-right (row 1, col. 2) cell of Figure 1, the results were satisfactory, as most of the returned results include a banana (it also returned some irrelevant results, such as an image of a person wearing a banana costume). However, it failed for a moderately complex query, [An unripe banana], as none of the returned images was an unripe (green) banana – the middle-right (row 2, col. 2) cell of Figure 1. It further failed for the more complex queries, including the one in Example 1.1, and was not able to find the image of interest in its results. – the bottom-right (row 3, col. 2) cell of Figure 1. As we shall further explain in § 6, we observed a consistent behavior for this approach across a wide range of queries.
This motivates us to take a step toward filling this gap in this paper by proposing a Monte Carlo randomized algorithm. In particular, we empower our algorithm with generative AI (GenAI) to generate a set of synthetic multi-modal tuples that represent the complex natural language query. Then, leveraging a collection of embedders, we conduct a series of NN-search using the generated tuples and aggregate the results to generate the final query outputs. We use our Monte Carlo method to develop Needle, our system for complex natural query answering on image datasets.
Unlike existing approaches, Needle is capable of handling natural language queries with different levels of complexity. To demonstrate this, let us consider our running example, Example 1.1 once again. To compare Needle with CLIP, we issued the same set of simple, moderate, and complex queries. The results are provided in the first column of Figure 1. For the simple object-finding task, the top-left (row 1, col 1) cell in Figure 1, we observed a similar (but slightly better) performance since all returned images satisfied the query, i.e., they included a banana. It also performed reasonably well for the moderately complex query, as a large portion of the output images were green bananas – the middle-left (row 2, col 1) cell in Figure 1. But perhaps, more interestingly, it could successfully find the corresponding image described in the complex query of Example 1.1 – the bottom-left (row 3, col 1) cell in Figure 1. Interestingly, the target image was returned as the top-1 search result (the top-left image in the bottom-left cell), demonstrating the ability of our system to answer these types of queries. We observed a consistent behavior for our system in our experiments across a large number of queries (§ 6).
Summary of contributions
This paper introduces a new approach to using GenAI for complex natural-language query answering on multi-modal data. Accordingly, it proposes Needle, a system for image data retrieval. In summary, our contributions are the following:
  • We introduce a novel Monte Carlo method for answering complex natural-language queries on multi-modal data (§ 3). We leverage Foundation models to generate synthetic tuples representing natural language queries. We then use a collection of vector representations (aka embeddings) that allows us to apply traditional k-NN techniques to retrieve related tuples. Our Monte Carlo algorithm aggregates the output generated for each (synthetic tuple, embedding) pair to obtain the final result.
  • We use our Monte Carlo method to develop Needle, our interactive system image data retrieval (§ 4).
  • We propose a tiling technique, Smart Tiling, inspired by quadtrees (Finkel and Bentley, 1974), to detect and index objects in a densely populated image during the preprocessing time and tile the image based on the density of each image region (§ 4.2).
  • We propose a method to aggregate the top-$k$ rankings generated by each of the (synthetic tuple, embedding) pairs into a single final ranking while accounting for the reliability of each embedder (§ 5.1). This method enables readily adding/removing any of the current and future embedders to Needle.
  • We propose a LOF-based (Breunig et al., 2000) approach for detecting the poorly generated synthetic tuples as anomalies and discarding them (§ 5.2).
  • We conduct comprehensive experiments on multiple datasets to demonstrate the effectiveness of Needle in object detection benchmarks (§ 6.2.1), and its ability to handle scenarios with targeted queries (§ 6.2.2). We also explore the effect of different hyperparameters on Needle’s performance (§ 6.3) and the effectiveness of built-in anomaly detection method (§ 6.4).
  • Lastly, we conduct a case study where human evaluators compare Needle and OpenAI’s CLIP results, with 70% of participants preferring Needle over CLIP for their queries (§ 6.5).
2. Preliminaries
Data Model: We consider the dataset $\mathcal{D}=\{t_{1},t_{2},\ldots,t_{n}\}$ as a set of $n$ multi-modal tuples (e.g., images) with no explicit values on specific attributes or any meta-data description about them. This model complies with real-world needs, where personal devices such as cell phones have advanced cameras, enabling a vast volume of multi-modal data collection and sharing by any user of such devices.
Query Model: The input query $\varphi$ is defined as a natural language description that corresponds to a subset of tuples $\mathcal{D}_{\varphi}\subseteq\mathcal{D}$, where $\mathcal{D}_{\varphi}=\{t^{\prime}_{1},t^{\prime}_{2},\ldots,t^{\prime}_{k}\}$ represents the target subset intended for retrieval.
3. The GenAI-powered Monte-Carlo Method
Nearest-neighbor search in traditional settings is a well-studied line of research where many algorithms and indices such as Voronoi diagrams (Kolahdouzan and Shahabi, 2004), tree data structures (Yianilos, 1993), and local sensitive hashing (Jafari et al., 2021; Huang et al., 2015) have been proposed for efficient exact and approximate query answering. In such settings, the data points are specified as points in a $d$ dimensional space, i.e., $t_{i}\in\mathbb{R}^{d}$, where the goal is to find the $k$ nearest neighbors to a query point $q\in\mathbb{R}^{d}$ based on a distance function $dist:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$. That is, $k\mbox{-min}_{t_{i}\in\mathcal{D}}dist(q,t_{i})$. Such approaches, however, are not directly applicable in our setting because of the following challenges.
  1. (1)
    Multi-modal tuples such as images are not represented as points in $\mathbb{R}^{d}$.
  2. (2)
    The query is in the form of a (possibly complex) natural language phrase.
Our key idea for resolving these challenges is to leverage recent advancements in Generative AI (GenAI) and develop a randomized Monte Carlo algorithm that enables applying traditional nearest-neighbor techniques in our setting.
3.1. Vector Representations
Addressing the first challenge is possible using the vector representation (aka embedding) of the tuples (Oquab et al., 2014; Kiela and Bottou, 2014). A vector representation $\mathcal{E}:dom(t)\rightarrow\mathbb{R}^{d_{\mathcal{E}}}$ is a transformation of the tuples to a high-dimensional vector of numeric values where the semantic similarity of tuples is proportional to the cosine similarity of their vector representations. Specifically, given a vector representation $\mathcal{E}^{\ell}$, we present $\mathcal{E}^{\ell}(t_{i})$ as $\vec{v}^{\ell}_{i}=\langle v_{1},v_{2},\cdots,v_{d_{\mathcal{E}^{\ell}}}\rangle$. The distance between two vector embeddings $\vec{v}^{\ell}_{i}$ and $\vec{v}^{\ell}_{j}$ is computed based on their cosine similarity as
$\delta^{\ell}_{i,j}=\delta(\vec{v}^{\ell}_{i},\vec{v}^{\ell}_{j})=1-\cos\left(\angle(\vec{v}^{\ell}_{i},\vec{v}^{\ell}_{j})\right)$
Let $\Delta(t_{i},t_{j})$ be the (unknown) semantic distance of the tuples $t_{i}$ and $t_{j}$.
A vector representation $\mathcal{E}^{o}$ is optimal if it captures the semantic similarity of tuples, i.e., $\delta^{o}_{i,j}\sim\Delta(t_{i},t_{j}),~{}\forall t_{i},t_{j}$. Formally,
(1) $\Delta(t_{i},t_{j})\geq\Delta(t_{i},t_{k})~{}\Leftrightarrow~{}\delta^{o}_{i,j}\geq\delta^{o}_{i,k}$
In practice, however, vector representations are learned using a universe of available multi-modal data. As a result, instead of guaranteeing Equation 1, the learned embeddings satisfy a (weaker) guarantee in expectation. That is, the expected distance between the embeddings of two tuples is proportional to their semantic distance, i.e., $\mathbb{E}\left[\delta^{\ell}_{i,j}\right]\sim\Delta(t_{i},t_{j}),~{}\forall t_{i},t_{j}$.
3.2. A Baseline Approach
The state-of-the-art resolution for the second challenge is to develop vector representations that are jointly trained on the multi-modal data and textual data. Specifically, CLIP (Radford et al., 2021) is an embedding trained jointly on (images, text) pairs. As a result, CLIP transforms text and image to the vector-representation same space. Subsequently, given a natural language query, one can compute its CLIP embedding and find its nearest neighbor images in the embedding space. This approach, however, suffers from two major issues:
First, the paired text with the images are usually simple sentences (e.g., [A photo of a dog]), missing to fully describe the rich information encoded in the image. Conversely, our goal is to answer complex natural language queries, describing a scene with multiple objects and their relationships (e.g., [A dog sitting by the fireplace, staring at the cat on the couch]). As a result, as we shall verify it in our experiments (§ 6), while such embeddings may perform reasonably on simple natural language queries their performance drops significantly on complex queries.
Second, as explained in Section 3.1, learned embeddings only preserve semantic similarities in expectation. As a result, relying solely on one vector representation for query answering may be misleading and inaccurate. Therefore, we instead develop a randomized Monte Carlo algorithm that utilized a set of vector representation transformations to develop more accurate query answering.
3.3. Query Transformation using GenAI
In traditional nearest neighbor problems both the dataset tuples and the query are represented as points in the same space. However, in our problem the query is form of a natural language phrase, while the dataset contains multi-modal tuples.
Let $\varphi$ be the given natural language query, and let $\{\vec{v}^{o}_{1},\cdots,\vec{v}^{o}_{n}\}$ be the vector representations for the tuples in $\mathcal{D}$. If we could specify the corresponding vector $\vec{q}^{o}(\varphi)$ for $\varphi$ in the $\mathcal{E}^{o}$ space, we could find the $k$-nearest neighbors of $\varphi$ as
(2) $k\mbox{-NN}(\varphi,\mathcal{D})=\underset{t_{i}\in\mathcal{D}}{k\mbox{-argmin}}~{}\delta\left(\vec{q}^{o}(\varphi),\vec{v}^{o}_{i}\right)$
As a result, if we could generate the synthetic multi-modal tuple $\mathbf{g}_{\varphi}$ for which $\mathcal{E}^{o}(\mathbf{g}_{\varphi})=\vec{q}^{o}(\varphi)$, we could use Equation 2 for query answering.
Fortunately, recent advancements in generative AI, particularly the emergence of the foundation models (e.g., DALL·E 3) based on guided-diffusion techniques (Zhang et al., 2023; Rombach et al., 2022; Ramesh et al., 2022), have enabled the generation of synthetic multi-modal tuples for complex prompts $p$ (natural language descriptions of the tuple to be generated). However, due to many reasons such as the randomized nature of the foundation models, the generated tuples [^1] Throughout this paper, we refer to these synthetic tuples as “guide tuples” that represent the natural language query in search space. We use these tuples to “guide” our system toward retrieving the most relevant results. may not be the optimal synthetic tuple $\mathbf{g}_{\varphi}$, but we can assume that the generated tuple is (semantically) in the vicinity of $\mathbf{g}_{\varphi}$.
Formally, we assume every tuple generated by the foundation model for the prompt $p=toPrompt(\varphi)$ lies in an $\varepsilon$ distance from $\mathbf{g}_{\varphi}$. Where $\mathbb{E}[\varepsilon]=0$. In other words for a sample tuple $\bar{\mathbf{g}}_{\varphi}$ we have,
$\mathbb{E}\left[\Delta(\mathbf{g}_{\varphi},\bar{\mathbf{g}}_{\varphi})\right]=0$ $\Rightarrow\mathbb{E}\left[\delta\left(\mathcal{E}^{o}(\mathbf{g}_{\varphi}),\mathcal{E}^{o}(\bar{\mathbf{g}}_{\varphi})\right)\right]=0$
$\Rightarrow\mathbb{E}\left[\mathcal{E}^{o}(\bar{\mathbf{g}}_{\varphi})\right]=\mathcal{E}^{o}(\mathbf{g}_{\varphi})$
The final result follows from the principle that if the expected value of the cosine similarity between two vectors is 1, then the vectors are aligned in expectation.
3.4. The Monte Carlo Algorithm
We are now ready to develop our algorithm: a Monte Carlo randomized algorithm based on repeated sampling.
Given a natural language query $\varphi$, and at least one foundation model $\mathcal{F}$, at the first step, we generate $m$ synthetic guide tuples $\{\bar{\mathbf{g}}_{1},\cdots,\bar{\mathbf{g}}_{m}\}$. As mentioned before, we consider each guide tuple $\bar{\mathbf{g}}_{i}$ as an unbiased sample taken from a distribution with mean $\mathbf{g}_{\varphi}$.
Instead of relying on the distance estimations one vector representation, we view each embedding as a sample from the space possible transformation functions. Hence, we use multiple embedders to minimize the semantic distance estimation error. Let $\{\mathcal{E}^{1},\cdots,\mathcal{E}^{l}\}$ be the set of learned vector representations, considered for transformation. For each embedder $\mathcal{E}^{\ell}$, we assume $\mathbb{E}\left[\delta^{\ell}_{i,j}\right]=\delta^{o}_{i,j},~{}\forall t_{i},t_{j}$.
Using the generated tuples $\{\bar{\mathbf{g}}_{1},\cdots,\bar{\mathbf{g}}_{m}\}$ and the embedders $\{\mathcal{E}^{1},\cdots,\mathcal{E}^{l}\}$, we estimate the distance of each tuple $t_{i}$ to the query $\varphi$ as
(3) $\bar{\delta}_{\varphi,i}=\frac{1}{m\,l}\sum_{j=1}^{m}\sum_{\ell=1}^{l}~{}\delta\left(\mathcal{E}^{\ell}(\bar{\mathbf{g}}_{j}),\vec{v}^{\ell}_{i}\right)$
Theorem 3.1.
Let $\delta_{\varphi,i}=\delta\left(\vec{q}(\varphi),\vec{v}^{o}_{i}\right)$, $\forall t_{i}\in\mathcal{D}$. Given a positive value $\gamma$,
$\Pr\left(\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\geq(1+\gamma)\right)\vee\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\leq(1-\gamma)\right)\right)\leq\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{3}}+\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{2}}$
Proof.
Replacing $\bar{\delta}_{\varphi,i}$ with the right-hand side of Equation 3, we get
$\Pr\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\geq(1+\gamma)\right)$ $=\Pr\left(\frac{\frac{1}{m\,l}\sum_{j=1}^{m}\sum_{\ell=1}^{l}~{}\delta\left(\mathcal{E}^{\ell}(\bar{\mathbf{g}}_{j}),\vec{v}^{\ell}_{i}\right)}{\delta_{\varphi,i}}\geq(1+\gamma)\right)$
$=\Pr\left(\sum_{j=1}^{m}\sum_{\ell=1}^{l}~{}\delta\left(\mathcal{E}^{\ell}(\bar{\mathbf{g}}_{j}),\vec{v}^{\ell}_{i}\right)\geq(1+\gamma)m\,l\,\delta_{\varphi,i}\right)$
Since $\mathbb{E}\left[\delta^{\ell}_{i,j}\right]=\delta^{o}_{i,j},~{}\forall t_{i},t_{j}$ and $\mathbb{E}\left[\mathcal{E}^{o}(\bar{\mathbf{g}}_{\varphi})\right]=\mathcal{E}^{o}(\mathbf{g}_{\varphi})$,
$\mathbb{E}\left[\delta\left(\mathcal{E}^{\ell}(\bar{\mathbf{g}}_{j}),\vec{v}^{\ell}_{i}\right)\right]=\delta_{\varphi,i}$
As a result,
$\mathbb{E}\left[\sum_{j=1}^{m}\sum_{\ell=1}^{l}~{}\delta\left(\mathcal{E}^{\ell}(\bar{\mathbf{g}}_{j}),\vec{v}^{\ell}_{i}\right)\right]=m\,l\,\delta_{\varphi,i}$
Next, applying Chernoff bound, we get
$\Pr\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\geq(1+\gamma)\right)\leq\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{3}}$
Similarly,
$\Pr\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\leq(1-\gamma)\right)\leq\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{2}}$
Therefore, applying the union bound, we get
$\Pr\left(\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\geq(1+\gamma)\right)\vee\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\leq(1-\gamma)\right)\right)\leq\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{3}}+\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{2}}$
Using the estimated distances, we identify the $k$-nearest neighbors of $\varphi$ as
$\overline{k\mbox{-NN}}(\varphi,\mathcal{D})=\underset{t_{i}\in\mathcal{D}}{k\mbox{-argmin}}~{}\bar{\delta}_{\varphi,i}$
4. System Development
Figure 2. Needle Architecture
Having discussed the theoretical foundation of our system, we can now explain its architecture and different components in this section. While our GenAI-powered Monte Carlo method is modality agnostic, we will explain the development of our system for Image Retrieval. The overall architecture of Needle is illustrated in Figure 2. The architecture can be divided into two main parts: components involved in Preprocessing [§ 4.1, § 4.2, § 4.3] and Inference [§ 4.4, § 4.5, § 4.6, § 4.7].
4.1. Preprocessing
The preprocessing phase involves a one-time operation to create vector representations, also known as embeddings, of the input dataset using a set of predefined embedders. Since images are typically regarded as being strongly associated with their representing objects(Zhao et al., 2019), embedding models may underperform when applied to images containing numerous small objects (Liu et al., 2021b).
To address this limitation, we employ a tiling strategy that divides the input image into multiple tiles based on the density of objects in each region (§ 4.2).
Each of these tiles is then processed by the embedders, and the resulting embeddings are stored on disk. This process can be executed in parallel for each embedder, with performance significantly enhanced by using multiple GPUs to handle tile embedding.
4.2. Smart Tiling
Figure 3. An Illustration of Smart Tiling on an image with lots of objects
Algorithm 1 Smart Tiling
1:image, d, objects_borders, min_size
2:image tiles
3:$(h,w)\leftarrow\text{size of }image$
4:tiles [$\leftarrow$[]← [ ]
5:function CountObjects(x, y, w, h)
6:    return number of unique objects in the region $(x,y,w,h)$
7:end function
8:
9:function Split(x, y, w, h)
10:    num_objects $\leftarrow$ CountObjects(x, y, w, h)
[…]
12:         Append((x, y, w, h)) to tiles
13:    else
14:         Compute half_w, half_h
15:         Calculate objs_left, objs_right, objs_top, objs_bottom
16:         if vertical split is preferred & aspect ratio is reasonable then
17:             Split(x, y, half_w, h) $\triangleright$ Split vertically
18:             Split(x + half_w, y, half_w, h)
19:         else
20:             Split(x, y, w, half_h) $\triangleright$ Split horizontally
21:             Split(x, y + half_h, w, half_h)
22:         end if
23:    end if
24:end function
25:
26:Split(0, 0, w, h)
27:return tiles
To capture all the details within an image, it is important to focus on densely populated regions. Therefore, we should split each image into different tiles based on the density. In addition, most embedders are trained on square images (He et al., 2016; Simonyan and Zisserman, 2014; Szegedy et al., 2017) and perform optimally when applied to such shapes (Shahnaz and Mollah, 2022; Azulay and Weiss, 2019), which means excessively tall or narrow tiles should be avoided. In order to achieve these two goals, inspired by (Finkel and Bentley, 1974), we propose our Smart Tiling procedure. We employ an edge detection technique (Canny Edge Detection (Canny, 1986)) to identify object boundaries within the image and recursively split the image to balance the distribution of objects across tiles as effectively as possible. The objective of this algorithm is to divide the image into smaller tiles, where each tile contains no more than a specified number of objects $d$, while maintaining a balanced aspect ratio.
The algorithm recursively splits the image by first checking if a region contains more objects than the allowed threshold $d$. If this condition is met, the region is split either vertically or horizontally. The algorithm calculates the object distribution for potential splits and selects the one that provides the most balanced distribution of objects across the two resulting tiles. To avoid generating disproportionate tiles, the algorithm also checks the aspect ratio and favors splits that maintain near-square dimensions. This recursive process continues until each tile satisfies the object limit constraint or it reaches the minimum acceptable width/height, at which point the tiles are finalized. Algorithm 1 presents the pseudocode for our Smart Tiling approach.
4.3. Vector Store
Once the raw image data is processed into vectors, it is indexed within the vector store. The vector store provides a maximum inner product search interface with low latency, which is crucial since users expect prompt results from the system. In our implementation, we employ hnswlib, which offers approximate maximum inner product search capabilities. This is consistent with the functionality provided by most vector stores.
4.4. Query Refinement
Having discussed the components involved in Preprocessing, we now explore the components used during Inference. After receiving the query in natural language text, Needle refines the input query into a suitable prompt. The first step is to remove punctuation and redundant characters from the query. Additionally, Needle supports more advanced query refinement techniques.
[…]
In addition, given that current Foundation Models have limitations in fully understanding some types of prompts (Liu et al., 2022; Feng et al., 2022; Lee et al., 2023; Liu et al., 2024), our system can employ various query refinement techniques (Liu et al., 2021a; Liu and Chilton, 2022; Wang et al., 2023; Witteveen and Andrews, 2022) to paraphrase and reformat the input query. These transformations help better align the generated query tuples with the original intent. These techniques help overcome prompt misinterpretations and improve the relevance of generated results.
4.5. Foundation Model and Guide Image Generation
After refining the query, We need to generate guide tuples representing the query using foundation models. Needle can use multiple (black-box) foundation models to generate guide images, with the number of guide images fine-tuned as a hyperparameter, denoted as $m$. Foundation models are capable of producing images with varying levels of quality and resolution. However, for the purposes of Needle, high-quality realistic images are unnecessary to achieve accurate search results. Lower-resolution images are sufficient, provided that they closely resemble the input query. In our experiments (§ 6.3), we will explore the impact of $m$, the size of the generated images, and the overall quality of the foundation model in the final results.
4.6. User Interface and Feedback Process
Figure 4. Needle User Interface
The current implementation of Needle offers a web-based user interface that allows users to search based on their queries (Figure 4). Users can interact with the system in two modes: Feedback mode (on) or No-Feedback mode (off). When Feedback mode is enabled, the guide tuples are presented to the user prior to the search process. Users can evaluate the quality and relevance of the generated guide tuples with respect to their query, which can assist Needle in retrieving more relevant images. In No-Feedback mode, Needle automatically identifies and discards outlier tuples from the generated images without user intervention (§ 5.2).
4.7. Embedders and Rankings
After generating the guide images, Needle utilizes predefined embedders to convert the guide tuples into vector representations and performs a k-NN lookup on the vector stores (§ 4.3). The score of each image is determined by the maximum score among all of its tiles (with the entire image also treated as a tile). Subsequently, the results are aggregated using our aggregation mechanism (§ 5.1) to produce a final ranking, which is then presented to the user.
5. Practical Optimization
In this section, we discuss our practical optimizations to improve the efficiency of our system.
First, we address the issue of aggregating results from multiple embedders into a unified ranking (§ 5.1), followed by a discussion on detecting anomalies in generated synthetic guide tuples (§ 5.2).
5.1. Rankings Aggregation and Embedder Trust Mechanism
Needle has been designed with scalability and flexibility in mind, ensuring it is not restricted to any specific embedder. Particularly considering the rapid advancements in vector representations, our system is designed to dynamically add, remove, or upgrade embedders. This will allow our system to enhance its performance as more advanced embedders become available.
Needle aggregates the results of the available embedders in real-time to achieve optimal outcomes. Furthermore, Needle monitors the performance of embedders for different queries, adjusting its trust in them based on their historical performance. The performance of embedders may vary depending on the type of image input. For instance, one embedder may excel at distinguishing different breeds of dogs but underperform in fruit classification tasks, while another may perform better in the latter. Needle addresses this variability by dynamically adjusting the reliability score of embedders for various query topics.
Consider a (predefined) set of topics $T$, and let the natural language query $\varphi$ belong to the topic $t=t(\varphi)\in T$ [^2] The topic $t$ of query $\varphi$ is identified using a topic classification mechanism or is part of the query input.. Let $\mathbf{G}=\{\bar{\mathbf{g}}_{1},\cdots,\bar{\mathbf{g}}_{m}\}$ be the set of guide images generated for $\varphi$ and consider the embedders $\mathbf{E}=\{\mathcal{E}^{1},\cdots,\mathcal{E}^{l}\}$. For a guide image $\bar{\mathbf{g}}_{j}$, let $R^{i}_{j}=\mathcal{R}^{i}(\varphi,\bar{\mathbf{g}}_{j})$ be the sorted list of $k$ nearest images in $\mathcal{D}$ to the guide image $\bar{\mathbf{g}}_{j}$ according to the embedder $\mathcal{E}^{i}$.
Let $R^{i}_{j}$[r] be the $r$-th element of $R^{i}_{j}$.
For each topic $t\in T$, we assign a weight $w_{i}^{t}\in[0,1]$ to each embedder $\mathcal{E}^{i}$, representing the reliability of the embedder for answering the queries within topic $t$. The weights are normalized so that for each topic $t$,
$\sum_{i=1}^{l}w_{i}^{t}=1.$
5.1.1. Aggregation
Given the ranked lists from all the embedders and the topic $t$ of the query $\varphi$, the objective is to aggregate these lists into a single final ranking $R_{\varphi}$ by incorporating the topic-specific reliability weights $w^{t}_{i}$.
Let $I_{h}\in\mathcal{D}$ be a tuple returned by at least one of the embedders. Let $rank(I_{h},R^{i}_{j})$ be the rank assigned to $I_{h}$ according to the list $R^{i}_{j}$. Let $\mathsf{S}:[k]\rightarrow[0,1]$ be a monotonically decreasing function that specifies the importance of each rank position in the top-$k$. In particular, while being agnostic to the choice of $S(.)$, following the literature (Cummins, 2012; Arampatzis and Robertson, 2011; Faria et al., 2010), we adopt the following position-importance function in Needle:
$\mathsf{S}(i)=\begin{cases}\frac{1}{i}&~{}i\leq k\\ 0&~{}\text{otherwise}\end{cases}$
Then, the aggregated score of $I_{h}$ is computed as:
(4) $score(I_{h})=\sum_{i:\mathcal{E}^{i}\in\mathbf{E}}w^{t}_{i}\cdot\mathsf{S}(rank(I_{h},R^{i}_{j}))$
The top-$k$ tuples based on Equation 4 are returned as the query output, denoted as $R_{\varphi}$.
5.1.2. Dynamic Reliability-weight Adjustment
Suppose after showing the top-$k$ results $R_{\varphi}$ to the user, we receive feedback from him/her, which is a subset of results marked as irrelevant to the query $\varphi$  [^3] We acknowledge that such feedback may not commonly be available after the system deployment. However, crowdsourced feedback on a small set of topic-based benchmarks can be employed to calculate these weights before deploying a new embedder to our system..
Let $\mathcal{T}=\{I^{\prime}_{1},I^{\prime}_{2},\dots,I^{\prime}_{s}\}$ be the set of irrelevant tuples where $\mathcal{T}\subseteq R_{\varphi}$
Using this feedback, we update the reliability weights of the embedders for each topic using the Multiplicative Weight Update Method (MWUM) (Arora et al., 2012). Let $t=t(\varphi)$ be the topic of the query $\varphi$. Then the partial loss of the embedder $\mathcal{E}^{i}$ for a topic $t$ based on $\varphi$ is computed as,
(5) $\Delta loss_{\varphi}(t,i)=\sum_{j:\bar{\mathbf{g}}_{j}\in\mathbf{G}}\sum_{I^{\prime}_{h}\in\mathcal{T}}\mathsf{S}(rank(I^{\prime}_{h},R^{i}_{j}))$
We only penalize those embedders that highly ranked the irrelevant results within the query topic $t$.
The topic-specific weights are updated multiplicatively:
$w_{i}^{t}(t+1)=w_{i}^{t}(t)\cdot(1-\eta\cdot\Delta loss_{\varphi}(t,i))$
where $\eta\in(0,1)$ is the learning rate controlling how much the trust in each embedder changes over time. Finally, the weights are normalized for topic $t$ to ensure $\sum_{i=1}^{l}w_{i}^{t}=1$.
5.2. Anomaly Detection in Query Images
In addition to rankings aggregation, Needle employs an outlier detection mechanism to identify potentially anomalous guide images based on their embeddings from multiple embedders. These anomalies might indicate data quality issues or images that are inherently different from the majority. This process combines the embeddings from all available embedders and uses a weighted approach to adjust for the reliability of each embedder.
We are given a set of $m$ guide images $\mathbf{G}=\{\bar{\mathbf{g}}_{1},\bar{\mathbf{g}}_{2},...,\bar{\mathbf{g}}_{m}\}$ and a set of $l$ embedders $\mathbf{E}=\{\mathcal{E}^{1},\mathcal{E}^{2},...,\mathcal{E}^{l}\}$. For each image $\bar{\mathbf{g}}_{j}$, the embedder $\mathcal{E}^{i}$ produces a high-dimensional embedding $\vec{v}^{i}_{j}=\mathcal{E}^{i}(\bar{\mathbf{g}}_{j})$. Thus, for each image $\bar{\mathbf{g}}_{j}$, we have $l$ corresponding embeddings from the embedders:
$\{\vec{v}^{1}_{j},\vec{v}^{2}_{j},\dots,\vec{v}^{l}_{j}\}.$
Our goal is to detect if any guide image $\bar{\mathbf{g}}_{j}$ is an anomaly by analyzing the embeddings across all embedders.
5.2.1. Dimensionality Reduction using UMAP (McInnes et al., 2018)
Given the high-dimensional nature of the embeddings, we first apply dimensionality reduction to project the embeddings into a lower-dimensional space. For each embedder $\mathcal{E}^{i}$, we reduce the dimensionality of the embeddings $\{\vec{v}^{i}_{1},\vec{v}^{i}_{2},\dots,\vec{v}^{i}_{m}\}$ to a lower dimension $d$ using UMAP (Uniform Manifold Approximation and Projection) (McInnes et al., 2018).
UMAP constructs a weighted $k$-nearest neighbors graph of the high-dimensional points, optimizes the layout in a low-dimensional space, and minimizes the cross-entropy between the two spaces. The reduced embeddings are given by:
$\vec{v^{\prime}}^{i}_{j}=\text{UMAP}(\vec{v}^{i}_{j})\in\mathbb{R}^{d}$
where $d\ll\text{dim}(\vec{v}^{i}_{j})$.
5.2.2. Outlier Detection using LOF (Breunig et al., 2000)
After dimensionality reduction, we apply the Local Outlier Factor (LOF)(Breunig et al., 2000) algorithm to identify outliers. LOF assigns an outlier score to each data point based on the density of its neighborhood. The LOF score for image $\bar{\mathbf{g}}_{j}$ with respect to embedder $\mathcal{E}^{i}$ is defined as:
$\text{LOF}^{i}(\bar{\mathbf{g}}_{j})=\text{LOF}(\vec{v^{\prime}}^{i}_{j})$
This score indicates how much of an outlier $\bar{\mathbf{g}}_{j}$ is in the context of the embeddings from embedder $\mathcal{E}^{i}$, with higher scores representing a greater likelihood of being an outlier.
5.2.3. Aggregation of Outlier Scores
The final outlier score of an guide image $\bar{\mathbf{g}}_{j}$ is computed by aggregating the individual scores, weighted with the reliability weight of each embedder:
$S(\bar{\mathbf{g}}_{j})=\sum_{i=1}^{l}w_{i}\cdot\text{LOF}^{i}(\bar{\mathbf{g}}_{j})$
5.2.4. Anomaly Detection Threshold
In the end, if the outlier score $S(\bar{\mathbf{g}}_{j})$ of an image $\bar{\mathbf{g}}_{j}$ exceeds a given threshold $\tau$, the image is flagged as an anomaly:
$S(\bar{\mathbf{g}}_{j})>\tau\quad\Rightarrow\quad\bar{\mathbf{g}}_{j}\text{ isan anomaly}$
The threshold $\tau$ is a hyper-parameter adjusted according to the sensitivity required by the user. Higher values of $\tau$ promote greater diversity in the retrieved images, potentially increasing recall but risking the inclusion of images that do not adequately represent the query. This may result in lower accuracy and decreased precision. In contrast, lower values of $\tau$ force Needle to adhere more closely to the query, leading to higher precision in the results but potentially missing some relevant images.
6. Experiments
This section presents results of our experiments designed to evaluate the effectiveness of our proposed system, Needle. We begin by outlining the experimental setup (§ 6.1) and proceed with a proof of concept (PoC) (§ 6.2), where we demonstrate Needle ’s ability to retrieve relevant images across different scenarios.
In the first scenario, we focus on object detection tasks, which may involve identifying either rare (single occurrence) or common objects within the dataset (§ 6.2.1). In the second scenario, the goal is to locate a specific image using its natural language description. We evaluate how accurately Needle can retrieve the target tuple based on the provided description.
Following the PoC, we explore the impact of various hyper-parameters on Needle’s performance (§ 6.3) and how can Needle perform well, even with a few low-resolution guide tuples. We then analyze the performance of the anomaly detection component embedded within Needle and will compare it with human evaluation results(§ 6.4). Finally, we present a case study in which human evaluators compare Needle ’s results with CLIP (§ 6.5).
6.1. Experimental Setup
6.1.1. Datasets
We have used Caltech256 (Griffin et al., 2022), LVIS (Gupta et al., 2019), and the Berkeley DeepDrive (BDD) (Yu et al., 2020) datasets for object detection. Caltech256 is a relatively simple benchmark that includes 256 categories of objects with 30,000 images. LVIS is a long-tailed benchmark with over 1,300 categories and 100,000 images. Due to the fact that some categories in LVIS have fewer than 10 images across the entire dataset, this benchmark is considered particularly challenging for object detection tasks. BDD is a large-scale, diverse dataset for autonomous driving, which contains over 100,000 video frames and images annotated for object detection, lane detection, and drivable area segmentation.
For the targeted queries experiments, we use two image captioning benchmarks: Nocaps (Agrawal et al., 2019) and SentiCap (Mathews et al., 2016). Nocaps is a large-scale image captioning dataset that extends the Open Images dataset (Kuznetsova et al., 2020), featuring captions for novel object categories and their relations to each other. SentiCap (Mathews et al., 2016), on the other hand, focuses on sentiment-driven image captions, containing images annotated with captions that convey either positive or negative sentiments, making it suitable for evaluating models on emotional and subjective image descriptions.
6.1.2. Evaluation Metric
In our experiments, we report Average Precision (AP) as the primary evaluation metric. AP is a widely used accuracy metric in information retrieval, as it places greater emphasis on retrieving relevant results earlier in the ranking compared to simpler metrics such as precision or F1-score. AP values range from 0 to 1, with an AP of 1 indicating perfect accuracy, meaning all relevant results appear before any irrelevant ones. For a given query, AP is calculated as the average of the individual precision scores at each possible recall threshold: $AP=(\sum_{i=1}^{R}P_{i})/R$, where $R$ is the total number of relevant results in the dataset, and $P_{i}$ represents the precision when considering the top $i$ relevant results. In addition to reporting the mean Average Precision (mAP) as the final performance indicator for each baseline, we also report Hit Rate for targeted queries. Since these queries are designed to find exactly one target image, Hit Rate measures whether the relevant image appears in the results, regardless of its rank. The mean Hit Rate (mHR) across all queries is also presented for each system.
6.1.3. Baselines
In our experiments, we use the latest and highest-performing version of CLIP, specifically $ViT$-$B/32$, as the baseline model (Radford et al., 2021). CLIP (Contrastive Language–Image Pretraining) is a powerful vision-language model developed by OpenAI, designed to learn visual concepts from natural language descriptions. It achieves this by jointly training an image encoder and a text encoder using a contrastive learning objective, aligning images and their corresponding textual descriptions in a shared embedding space. CLIP has demonstrated strong zero-shot performance across a wide range of visual tasks, making it a robust baseline for image retrieval tasks.
In all experiments, we will report AP for all queries in the dataset and their “hard subset”. For each dataset, we define the hard subset as the set of tuples for which CLIP AP is less than 0.5.
6.1.4. Environmental Setup
All experiments were conducted on a local server equipped with 128 GB of memory, 32 CPU cores, and two NVIDIA GeForce RTX 2080 Ti GPUs. Four different Foundation Models were utilized throughout the experiments. runwayml-v1.5 and sdxl-turbo-v1.0 were hosted locally on the aforementioned server, while DALL·E 3 and Flux-Schnell were accessed via paid APIs.
6.2. Proof of Concepts
6.2.1. Object Detection Experiments
Table 1. Performance (mAP) Comparison Between CLIP and Needle on Object Detection Benchmark Datasets
Dataset (#hard/#all) All Queries Hard Subset
CLIP Needle CLIP Needle
BDD (2/11) 0.76 0.83 0.01 0.24
Caltech256 (10/256) 0.94 0.95 0.26 0.59
LVIS (1049/1203) 0.16 0.31 0.08 0.23
Average 0.62 0.70 0.12 0.35
((a)) LVIS
((b)) Caltech256
((c)) BDD
Figure 5. Distribution of Average Precision for different queries for CLIP and Needle, split into hard and easy subsets.
In these experiments, we demonstrate the effectiveness of Needle in basic object detection tasks. We utilize state-of-the-art benchmark datasets, LVIS (Gupta et al., 2019), Caltech256 (Griffin et al., 2022), and BDD (Yu et al., 2020), for this evaluation. For each query, the number of retrieved images is set to 60 ($k=60$), and we aim to identify 10 relevant tuples within the results ($R=10$). Two local Stable Diffusion models are employed as Foundation Models: runwayml-v1.5 and sdxl-turbo-v1.0 with guide image resolution of $768\times 768$. For each model, two guide images are generated, resulting in a total of four guide images per query ($m=4$). Table 1 presents a comparative analysis of performance between CLIP and Needle across the different datasets. Notably, Needle consistently demonstrates superior performance across all categories, in particular, for the hard subsets, Needle achieves around 200% improvement in AP compared to CLIP on average, highlighting its effectiveness in addressing more challenging object detection scenarios. Figure 5 shows the AP distribution of CLIP and Needle for different datasets. Needle demonstrates significant improvements in AP, even for queries where CLIP initially performs very poorly, with AP close to zero.
It is important to note that all generated guide images are produced using local Stable Diffusion models, which generate lower quality images than commercial alternatives. This result highlights that Needle imposes much looser constraints on the quality of guide tuples than humans do. Even when the guide images are distorted or lack realism, Needle can robustly retrieve relevant images using these guide tuples.
6.2.2. Targeted Queries
Figure 6. Illustration of different specific queries, the generated guide images and the comparison of CLIP and Needle results
Table 2. Performance (mAP, mHR) Comparison Between CLIP and Needle on Targeted Queries Benchmark Datasets
Dataset (#queries) mAP mHR
CLIP Needle CLIP Needle
Nocaps (100) 0.11 0.37 0.55 0.73
SentiCap (100) 0.24 0.34 0.71 0.77
Average 0.18 0.36 0.63 0.75
Needle excels at handling targeted queries and identifying rare tuples. We demonstrate this by running two image captioning benchmarks, Nocaps (Agrawal et al., 2019), and SentiCap (Mathews et al., 2016), using both CLIP and Needle. For each experiment, the systems are provided with a query that describes a specific image in the dataset, with the search engines returning 10 images. For this task, we enhance Needle with two more powerful and precise Foundation Models: DALL·E 3  [^4] https://openai.com/index/dall-e-3/ and Flux-Schnell  [^5] https://replicate.com/black-forest-labs/flux-schnell. These models are selected due to the need for guide images that accurately represent the detailed input queries, as these queries include significant complexity and detail.
In addition to reporting mean Average Precision (mAP), we also evaluate the mean Hit Rate (mHR). Hit Rate is particularly useful in cases where the number of retrieved images is small, and the focus is on finding the correct image rather than its specific ranking.
As shown in Table 2, Needle significantly outperforms CLIP in locating the target image. Needle successfully retrieves the correct image in over 75% of queries within the top 10 results, marking around 20% improvement over CLIP’s performance in mHR and 100% improvement in mAP. Figure 6 presents four different queries from the Nocaps (Agrawal et al., 2019) dataset, along with the generated guide images for Needle and the top-1 results from both CLIP and Needle. As illustrated, CLIP frequently overlooks key aspects of the input prompt and focuses on a single keyword rather than capturing the full intent of the query. In contrast, Needle effectively captures all inherent complexity of query within its guide tuples and identifies the target image using only two guide images.
6.3. Parameter Tuning
((a)) Runwayml (Caltech256)
((b)) sdxl-turbo (Caltech256)
((c)) Runwayml (LVIS)
((d)) sdxl-turbo (LVIS)
Figure 7. Heatmaps of Needle mAP for two different datasets (Caltech256 and LVIS) using two local Foundation Models (Runwayml and sdxl-turbo) while altering number of Generated Images $m$ and Generated Image Size
In this section, we investigate the effect of varying the number of generated guide images ($m=\{1,2,4\}$), the image resolution ($\{512\times 512,768\times 768,1024\times 1024\}$), and the performance of different foundation models (runwayml, sdxl-turbo) on Needle’s effectiveness. The experiments are conducted on both the Caltech256 and LVIS datasets. Figure 7 illustrates the mAP values for different combinations of these parameters. As a general trend, increasing the number of images and the resolution improves mAP performance. However, there are deeper insights to consider: first, sdxl-turbo exhibits more consistent and slightly better performance across different image sizes compared to runwayml; second, while increasing the number of generated images $m$ improves results, the improvement diminishes and stabilizes after a certain point. A similar trend is observed for image sizes. These results suggest that Needle can maintain high performance even with a limited number of generated images and lower resolution inputs.
6.4. Feedback Mode
((a)) LVIS
((b)) BDD
Figure 8. Comparison of AP distribution for Needle on Feedback mode on and off
CLIPNeedleBothNeither002020202040404040606060608080808010010010010023.2323.2323.2323.2352.5252.5252.5252.5214.1514.1514.1514.1510.110.110.110.1Percentage (%)
Figure 9. Case study results
Table 3. Comparison of Needle’s performance (mAP) for different datasets with Feedback mode enabled (ON) and built-in Anomaly detection (OFF)
Dataset (#hard/#all) All Queries Hard Subset
OFF ON OFF ON
BDD (2/11) 0.82 0.84 0.22 0.25
LVIS (84/100) 0.28 0.30 0.19 0.21
Average 0.55 0.57 0.21 0.23
In this subsection, we investigate the impact of enabling the feedback procedure in Needle on mAP. A human evaluator was employed to verify the quality and alignment of the generated images with the input query prior to retrieving results. For evaluation, we used LVIS and BDD datasets, generating four tuples per query ($m=4$) with an image size of $768\times 768$, utilizing the runwayml foundation model. We compare the mAP of Needle’s built-in anomaly detection method (§ 5.2) to the performance when human feedback is incorporated.
Figure 8 presents the AP distribution for both baselines. As shown, there is a slight improvement when feedback mode is enabled, primarily raising queries with $AP\approx 0.4$ in LVIS and $AP\approx 0.8$ in BDD toward higher APs. This indicates that: (a) Needle’s built-in anomaly detection method performs effectively in practice, achieving results comparable to a human evaluator, and (b) human intervention is most beneficial for queries that already have a moderate AP, whereas queries for which foundation models fail to generate suitable images do not significantly improve with feedback alone.
6.5. Case Study
To further validate Needle’s practical applicability, we conducted a case study with human evaluators. We utilized the LVIS dataset as our primary image repository, as it contains over 100k images spanning various categories, ensuring that any randomized query would likely have related images. This made LVIS an ideal choice for our case study.
Each human evaluator was asked to write five queries of their choice, with varying levels of complexity. For each query, we retrieved 20 relevant images from both CLIP and Needle, displaying the results side by side with randomized orientations for each query. Evaluators were then asked to select the more relevant results, choosing from four options: ”Left is better,” ”Right is better,” ”Both are good enough,” and ”Neither is good enough.”, then we recorded the engine behind that option as preferred engine.
A total of 20 human evaluators participated in the study, contributing 100 queries. The queries ranged from simple ones like ”Celery stew” to more complex descriptions such as ”A person walking on the sidewalk of a river in Fall at Chicago.”, Figure 9 presents the results of this user study. If we consider only the queries where one engine is preferred over the other, Needle is chosen as the better engine in almost 70% of cases.
7. Related Work
Our paper relates to work in the following areas:
Multi-modal Data Retrieval: Multi-modal data retrieval, which integrates various data types (e.g., text, images, video, and audio), has attracted significant attention. Early methods focused on shared representations by mapping modalities into a common latent space using canonical correlation analysis (CCA) (Rasiwasia et al., 2010; Wang et al., 2015b) and kernel-based techniques (Jia et al., 2019). Recent approaches leverage deep learning models, including cross-modal encoders (Wang et al., 2015a; Liu et al., 2017) and joint embedding networks (Balaneshin-kordan and Kotov, 2018; Rubio et al., 2017). Pre-trained transformers like CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) highlight the effectiveness of combining vision and language models for cross-modal tasks.
Advanced techniques, such as attention mechanisms (Yang et al., 2021; Tang et al., 2023) and contrastive learning (Pham et al., 2024), further enhance retrieval by capturing fine-grained relationships. Feedback mechanisms (Moll et al., 2023; Zhou and Huang, 2003) and user preferences (Zhang and Zhang, 2010; Chen et al., 2020) improve multi-modal search results. Research also extends to other modalities, such as text-video retrieval (Li et al., 2019; Markatopoulou et al., 2017; Torabi et al., 2016; Mithun et al., 2018; Miech et al., 2019; Luo et al., 2021) and text-audio retrieval (Xin et al., 2023; Chechik et al., 2008; Koepke et al., 2022). For a detailed survey of cross-modal retrieval methods and future directions, see (Li et al., 2023).
GenAI in Data Management: The intersection of Generative AI (GenAI) and data management has emerged as a promising area, driving innovations in how data is generated, managed, and utilized. GenAI models, such as GPT (Achiam et al., 2023) and Stable Diffusion (Rombach et al., 2022), have been employed for synthetic data generation to augment datasets (Tanaka and Aranha, 2019; Shahbazi et al., 2024), improve model training (Le et al., 2017; Tremblay et al., 2018), and address data scarcity or imbalance issues (Erfanian et al., 2024). In data cleaning and integration, these models assist with tasks like imputation (Villaizán-Vallelado et al., 2024), and anomaly detection by generating plausible data patterns (Ye et al., 2024). Recent works also explore the role of GenAI in automating query generation (Trummer, 2022, 2023; Li and Xie, 2024) and data summarization (Dar et al., 2024), enhancing the efficiency of database systems and analytics platforms.
Natural Language Query Processing: The evolution of natural language to structured query translation has advanced from rule-based systems to sophisticated large language models (LLMs). Early methods used template matching and predefined rules (Androutsopoulos et al., 1995; Li and Jagadish, 2014), which were inflexible and required frequent manual updates. Machine learning approaches introduced more adaptability by learning query patterns from annotated datasets (Ma et al., 2020; Zhang et al., 2019), though they struggled with complex language constructs.
Deep learning models, especially transformers like BERT (Devlin, 2018) and GPT (Brown, 2020), enabled semantic understanding and laid the foundation for text-to-SQL systems (Li et al., 2024). Modern LLMs, such as PaLM (Chowdhery et al., 2023) and GPT-4 (Achiam et al., 2023), generate context-aware queries across multiple domains with minimal supervision. Retrieval-augmented generation (RAG) further enhances query precision by leveraging external knowledge (Lewis et al., 2020). Ongoing efforts focus on fine-tuning LLMs to improve accuracy and efficiency in converting natural language to structured queries (Trummer, 2022, 2023; Li and Xie, 2024).
8. Discussions and Limitations
8.1. Reliance on Existing Models
Needle extensively relies on Foundation Models for guide tuple generation and on embedders for generating vector representations of images. Both these components have their own capabilities, limitations, and inherent biases. This reliance presents both strengths and weaknesses in Needle’s architecture. The strength lies in the ability to easily improve performance by upgrading the internal Foundation Models and embedders. However, the limitations of these models also become limitations for Needle. To mitigate this, we designed Needle to support multiple Foundation Models and embedders, enabling it to draw from a broader range of knowledge and reduce the impact of individual model limitations.
8.2. Efficiency
We assess the efficiency of Needle across two phases: Preprocessing and Inference.
Preprocessing: In the preprocessing phase, two highly parallelizable tasks are performed: tiling and embedding creation. Tiling is a fast operation, but it can linearly increase the time required for embedding creation. To improve time efficiency, one can increase the minimum tile width and height, thereby reducing the number of tiles. The time required for embedding creation can be significantly improved by utilizing multiple GPUs. Since the preprocessing phase is a one-time operation that can be easily parallelized and fine-tuned, it does not present a major concern for Needle’s overall performance.
Inference: Needle’s inference consists of three main sequential operations: (a) generating guide tuples, (b) k-NN lookup, and (c) ranking aggregation. For an input dataset of size $n$, retrieving $k(1\leq k\leq n)$ results with constants $m$ (number of guide images) and $l$ (number of embedders), (a) is a constant-time operation with respect to $n$ and $k$. Step (b)’s time complexity depends on the underlying vector store algorithm, and for HNSW data structure (Malkov and Yashunin, 2018), is $O(k\log(n))$. Step (c) has a time complexity of $O(k)$ using Fagin’s Algorithm (Fagin et al., 2001).
In our experiments, where the number of tuples (tiles) is at most 400,000 for any k, (b) k-NN lookup and (c) the rankings aggregation take less than one second using current approximate vector stores. The bottleneck of the entire process lies in (a), generating guide tuples. The good news is that this operation is independent of the dataset size $n$ and number of images to retrieve $k$, and only relies on the number of guide images $m$. Moreover, the generation process can run in parallel, significantly improving performance.
To summarize, for small values of $n$, Needle’s inference time may be slower than other methods due to the constant time required for guide image generation. However, as $n$ increases, the effect of step (a) diminishes, and Needle’s efficiency becomes comparable to other methods.
8.3. Extensions to Other Modalities
The core idea behind Needle is adaptable to various modalities beyond images, including audio, video, and others. However, due to the current limitations in publicly available foundation models for audio and video generation, our focus in this paper has been restricted to the image modality. As foundation models for these additional modalities continue to improve and become more accessible, we plan to extend Needle to support them. This will allow Needle to operate across multiple data types, unlocking further potential for handling complex, multi-modal queries in the near future.
9. Conclusion
In this paper, we took a step towards addressing the challenge of answering complex natural language queries in multi-modal data settings by proposing a Generative-AI Powered Monte Carlo method. Our proposed method utilizes foundation models to generate synthetic guide tuples that capture the complexities of the input query; then, we leverage these guide tuples to represent the natural language query in multi-modal data space and apply traditional nearest neighbor methods.
Following this method, We implemented Needle for image retrieval. We proposed different solutions for practical challenges in our system’s development. Our comprehensive experiments demonstrated that our system, Needle, significantly outperforms the state-of-the-art techniques in image retrieval.
While our proposed technique is not limited to image datasets, its development for other modalities, such as audio and video, requires foundation models that are currently not publicly available. Upon the availability of such models, we aim to extend the scope of Needle to a wider range of modalities in our future work.

### v2

@@
-Needle: A Generative-AI Powered Monte Carlo Method for Answering Complex Natural Language Queries on Multi-modal Data
-Abstract.
-Multi-modal data, such as image data sets, often miss the detailed descriptions that properly capture the rich information encoded in them.
-This makes answering complex natural language queries a major challenge in these domains.
-In particular, unlike the traditional nearest-neighbor search, where the tuples and the query are modeled as points in a data cube, the query and the tuples are of different natures, making the traditional query answering solutions not directly applicable for such settings.
-Existing literature addresses this challenge for image data through vector representations jointly trained on natural language and images.
+Needle: A Generative AI-Powered Multi-modal Database for Answering Complex Natural Language Queries
+Abstract
+Multi-modal datasets, like those involving images, often miss the detailed descriptions that properly capture the rich information encoded in each item.
+This makes answering complex natural language queries a major challenge in this domain.
+In particular, unlike the traditional nearest neighbor search, where the tuples and the query are represented as points in a single metric space, these settings involve queries and tuples embedded in fundamentally different spaces, making the traditional query answering methods inapplicable.
+Existing literature addresses this challenge for image datasets through vector representations jointly trained on natural language and images.
 This technique, however, underperforms for complex queries due to various reasons.
-This paper takes a step towards addressing this challenge by introducing a Generative-AI (GenAI) powered Monte Carlo method that utilizes foundation models to generate synthetic samples that capture the complexity of the natural language query and transform it to the same space of the multi-modal data.
-Following this method, we develop a system for image data retrieval and propose practical solutions that enable leveraging future advancements in GenAI and vector representations for improving our system’s performance.
-Our comprehensive experiments on various benchmark datasets verify that our system significantly outperforms state-of-the-art techniques.
-Image Databases; Generative AI; Foundation Models;
-1. Introduction
-a Needle in a haystack: something that is impossible or extremely difficult to find, especially because the area you have to search is too large.”
-Cambridge Dictionary
-Figure 1.
-The query results for the queries: (a) [a banana] (row 1), (b) [an unripe banana] (row 2), and (c) [a banana gazing at its reflection in a mirror] (row 3), using Needle (first column) v.s.
-CLIP (second column).
-Multi-modal data, including image datasets, pose new challenges for data management systems.
-Unlike traditional databases, where tuple attributes are explicitly specified, multi-modal data often miss proper descriptions that capture the rich information encoded in them.
+This paper takes a step towards addressing this challenge by introducing a Generative-based Monte Carlo method that utilizes foundation models to generate synthetic samples that capture the complexity of the natural language query and represent it in the same metric space as the multi-modal data.
+Following this method, we propose Needle, a database for image data retrieval.
+Instead of relying on contrastive learning or metadata-searching approaches, our system is based on synthetic data generation to capture the complexities of natural language queries.
+Our system is open-source and ready for deployment, designed to be easily adopted by researchers and developers.
+The comprehensive experiments on various benchmark datasets verify that this system significantly outperforms state-of-the-art text-to-image retrieval methods in the literature.
+Any foundation model and embedder can be easily integrated into Needle to improve the performance, piggybacking on the advancements in these technologies.
+1 Introduction
+Multi-modal datasets, like images, pose new challenges for data management systems.
+Unlike traditional databases, where tuple attributes are explicitly specified, these datasets often miss proper descriptions that capture the rich information encoded in each item.
 As a result, traditional query-answering approaches are not directly applicable in these settings.
-On the other hand, advanced personal devices such as cell phones have enabled the collection of large volumes of multi-modal data, even at the individual user level.
-Due to the wealth of information “hidden” in each multi-modal tuple, the users may compose complex natural language queries, describing the tuple(s) they are interested in.
+Meanwhile, modern personal devices like smartphones have made it possible to collect large amounts of multi-modal data, even from individual users.
+Due to the wealth of information “hidden” in each multi-modal tuple, the users may compose complex natural language queries, searching for the tuple(s) they are interested in.
 To further motivate this, let us consider the following running example.
-Example 1.1.
-A photo enthusiast has collected a large pool of images over the past several years on her private data repository.
+A photo enthusiast has collected a large pool of images over the past several years in her private data repository.
 She wants to retrieve a specific photo she has in mind.
 She describes the picture as [A banana gazing at its reflection in a mirror].
-Traditional nearest-neighbor (NN-) search algorithms are not directly applicable here since those consider the database tuples and the query as points in a data cube.
-In contrast, in Example 1.1, (i) each tuple is an image, and (ii) the query is in the form of an unstructured, complex natural language statement.
-The state-of-the-art approach for addressing the need for answering natural language queries on image datasets is CLIP (Radford et al., 2021), a vector representation (aka embedding) jointly trained on several hundred million (image, text) pairs.
-Using CLIP as the embedder, one can transform the query and the dataset images to vector representation in the same space and apply NN-search to answer the query.
-This approach works satisfactorily for “simple” queries, such as simple object detection tasks.
-To see a concrete example, following with our running example (Example 1.1), we used a benchmark image dataset (LVIS (Gupta et al., 2019)) and issued the query [A banana].
+[^1] See Figure 3 for more examples.
+Figure 1: The query results for (a) [a banana] (row 1), (b) [an unripe banana] (row 2), (c) [a banana gazing at its reflection in a mirror] (row 3), using Needle vs. CLIP [53].
+Traditional Nearest Neighbor (NN) search algorithms are not directly applicable here since they consider the dataset tuples and the query as points in the same data space.
+In contrast, in Example 1, (i) each tuple is an image, and (ii) the query is in the form of an unstructured, complex natural language statement.
+[…]
+As a concrete example, continuing with our running example (Example 1), we use CLIP [53], OpenAI’s well-known jointly-trained embedding model, to answer the natural language query [A banana] on the benchmark image dataset LVIS [22].
 As reflected in the top-right (row 1, col.
-2) cell of Figure 1, the results were satisfactory, as most of the returned results include a banana (it also returned some irrelevant results, such as an image of a person wearing a banana costume).
-However, it failed for a moderately complex query, [An unripe banana], as none of the returned images was an unripe (green) banana – the middle-right (row 2, col.
+2) cell of Figure 1, the results are satisfactory, as most of the returned images include a banana.
+However, it fails for a moderately complex query, [An unripe banana], as none of the returned images shows an unripe (green) banana – the middle-right (row 2, col.
 2) cell of Figure 1.
-It further failed for the more complex queries, including the one in Example 1.1, and was not able to find the image of interest in its results.
-– the bottom-right (row 3, col.
-2) cell of Figure 1.
-As we shall further explain in § 6, we observed a consistent behavior for this approach across a wide range of queries.
-This motivates us to take a step toward filling this gap in this paper by proposing a Monte Carlo randomized algorithm.
+It also fails on more complex queries, such as the one in Example 1, and does not retrieve the target image, the one shown in the bottom-right cell (row 3, col.
+2) of Figure 1.
+As we further explain in § 5, we observe a similar behavior to this example across a wide range of queries.
+This motivates us to take a step toward filling this gap by proposing a Monte Carlo randomized algorithm.
 In particular, we empower our algorithm with generative AI (GenAI) to generate a set of synthetic multi-modal tuples that represent the complex natural language query.
-Then, leveraging a collection of embedders, we conduct a series of NN-search using the generated tuples and aggregate the results to generate the final query outputs.
-We use our Monte Carlo method to develop Needle, our system for complex natural query answering on image datasets.
-Unlike existing approaches, Needle is capable of handling natural language queries with different levels of complexity.
-To demonstrate this, let us consider our running example, Example 1.1 once again.
-To compare Needle with CLIP, we issued the same set of simple, moderate, and complex queries.
+Subsequently, leveraging a collection of embedders, we conduct a series of NN searches using the generated tuples and aggregate the results to generate the final query outputs.
+We use our method to develop Needle  [^2] The name Needle is inspired by the common expression ”finding Needle in the haystack,” reflecting the challenge of pinpointing relevant information in large multimodal spaces., our deployment-ready open-source system333 Needle Github: https://github.com/UIC-InDeXLab/Needle/tree/main/ for complex natural query answering on image datasets.
+Unlike existing approaches, Needle can handle natural language queries with different levels of complexity.
+To demonstrate this, let us consider our running example, Example 1 once again.
+We issued the same set of simple, moderate, and complex queries to Needle.
 The results are provided in the first column of Figure 1.
@@
 It also performed reasonably well for the moderately complex query, as a large portion of the output images were green bananas – the middle-left (row 2, col 1) cell in Figure 1.
-But perhaps, more interestingly, it could successfully find the corresponding image described in the complex query of Example 1.1 – the bottom-left (row 3, col 1) cell in Figure 1.
+But perhaps, more interestingly, it could successfully find the image-of-interest described in the complex query of Example 1 – the bottom-left (row 3, col 1) cell in Figure 1.
 Interestingly, the target image was returned as the top-1 search result (the top-left image in the bottom-left cell), demonstrating the ability of our system to answer these types of queries.
-We observed a consistent behavior for our system in our experiments across a large number of queries (§ 6).
-Summary of contributions
-This paper introduces a new approach to using GenAI for complex natural-language query answering on multi-modal data.
-Accordingly, it proposes Needle, a system for image data retrieval.
+We observed a consistent behavior for our system in experiments across a large number of different queries (§ 5).
+Summary of contributions: This paper introduces a new approach to using GenAI for complex natural-language query answering on multi-modal data.
+We propose Needle, a system for image data retrieval.
 In summary, our contributions are the following:
-  • We introduce a novel Monte Carlo method for answering complex natural-language queries on multi-modal data (§ 3).
-We leverage Foundation models to generate synthetic tuples representing natural language queries.
+  • We introduce a novel Monte Carlo method for answering complex natural-language queries on multi-modal data.
+We leverage Foundation models to generate synthetic tuples representing the queries.
 We then use a collection of vector representations (aka embeddings) that allows us to apply traditional k-NN techniques to retrieve related tuples.
 Our Monte Carlo algorithm aggregates the output generated for each (synthetic tuple, embedding) pair to obtain the final result.
-  • We use our Monte Carlo method to develop Needle, our interactive system image data retrieval (§ 4).
-  • We propose a tiling technique, Smart Tiling, inspired by quadtrees (Finkel and Bentley, 1974), to detect and index objects in a densely populated image during the preprocessing time and tile the image based on the density of each image region (§ 4.2).
-  • We propose a method to aggregate the top-$k$ rankings generated by each of the (synthetic tuple, embedding) pairs into a single final ranking while accounting for the reliability of each embedder (§ 5.1).
-This method enables readily adding/removing any of the current and future embedders to Needle.
-  • We propose a LOF-based (Breunig et al., 2000) approach for detecting the poorly generated synthetic tuples as anomalies and discarding them (§ 5.2).
-  • We conduct comprehensive experiments on multiple datasets to demonstrate the effectiveness of Needle in object detection benchmarks (§ 6.2.1), and its ability to handle scenarios with targeted queries (§ 6.2.2).
-We also explore the effect of different hyperparameters on Needle’s performance (§ 6.3) and the effectiveness of built-in anomaly detection method (§ 6.4).
-  • Lastly, we conduct a case study where human evaluators compare Needle and OpenAI’s CLIP results, with 70% of participants preferring Needle over CLIP for their queries (§ 6.5).
-2. Preliminaries
-Data Model: We consider the dataset $\mathcal{D}=\{t_{1},t_{2},\ldots,t_{n}\}$ as a set of $n$ multi-modal tuples (e.g., images) with no explicit values on specific attributes or any meta-data description about them.
+(§ 3)
+  • We present Needle, a concrete, interactive system built upon our Monte Carlo method.
+Developed with production readiness in mind, Needle emphasizes ease of deployment, efficiency, robustness, and user-friendly interaction via a command-line interface (§ 4, Appendix D).
+  • We propose and integrate several practical optimizations designed to significantly improve Needle’s performance and efficiency, including a dynamic embedder trust mechanism, outlier detection for generated data, caching, and a query complexity classifier (§ 3.3).
+  • We conduct a comprehensive experimental evaluation demonstrating Needle’s effectiveness across diverse benchmarks (including object detection and complex natural language queries) compared to state-of-the-art baselines, supported by hyperparameter analysis and a human study indicating strong user preference for Needle (§ 5.2.1, § 5.2.2, Appendix E.4, Appendix E.5).
+2 Related Work
+Our paper relates to the following literature:
+Multi-modal Data Retrieval: Multi-modal data retrieval (aka cross-modal retrieval), integrating various data types, has recently attracted significant attention.
+Early methods focused on shared representations by mapping modalities into a common latent space using canonical correlation analysis (CCA) [55, 70] and kernel-based techniques [26].
+Recent approaches leverage deep learning models, including cross-modal encoders [69, 39] and joint embedding networks [6, 58].
+Advanced techniques, such as attention mechanisms [76, 62] and contrastive learning [52], further enhance retrieval by capturing fine-grained relationships.
+CLIP [53] and ALIGN [25] are widely adopted methods for cross-modal retrieval, both of which are based on contrastive learning approaches.
+Feedback mechanisms [48, 84] and user preferences [81, 11] improve multi-modal search results.
+In addition, there exists practical approaches to enrich the semantic metadata of images using image tagging [20, 83, 35] and object detection techniques [85].
+Research also extends to other modalities, such as text-video retrieval [36, 43, 64, 47, 46, 40] and text-audio retrieval [74, 10, 28].
+For a detailed survey of cross-modal retrieval methods and future directions, see [33, 71].
+To the best of our knowledge, this paper is the first to introduce the idea of utilizing GenAI to synthetically generate samples that capture natural language complexities and proposes a GenAI-powered Monte Carlo method for multi-modal data retrieval.
+GenAI in Data Management: The intersection of Generative AI (GenAI) and data management has emerged as a promising area, driving innovations in how data is generated, managed, and utilized.
+GenAI models, such as GPT [1] and Stable Diffusion [57], have been employed for synthetic data generation to augment datasets [61, 59], improve model training [30, 65], and address data scarcity or imbalance issues [16].
+In data cleaning and integration, these models assist with tasks like imputation [68], and anomaly detection by generating plausible data patterns [77].
+Recent works also explore the role of GenAI in automating query generation [66, 67, 37] and data summarization [14], enhancing the efficiency of database systems and analytics platforms.
+Natural Language Query Processing: The evolution of natural language to structured query translation has advanced from rule-based systems to sophisticated large language models (LLMs).
+Early methods used template matching and predefined rules [3, 32], which were inflexible and required frequent manual updates.
+Machine learning approaches introduced more adaptability by learning query patterns from annotated datasets [41, 82], though they struggled with complex language constructs.
+Deep learning models, especially transformers like BERT [15] and GPT [9], enabled semantic understanding and laid the foundation for text-to-SQL systems [34].
+Modern LLMs, such as PaLM [12] and GPT-4 [1], generate context-aware queries across multiple domains with minimal supervision.
+Retrieval-augmented generation (RAG) further enhances query precision by leveraging external knowledge [31].
+Ongoing efforts focus on fine-tuning LLMs to improve accuracy and efficiency in converting natural language to structured queries [66, 67, 37].
+3 The GenAI-powered Monte-Carlo Method
+We consider the dataset $\mathcal{D}=\{t_{1},t_{2},\ldots,t_{n}\}$ as a set of $n$ multi-modal tuples with no explicit values on specific attributes or meta-data description about them.
 This model complies with real-world needs, where personal devices such as cell phones have advanced cameras, enabling a vast volume of multi-modal data collection and sharing by any user of such devices.
-Query Model: The input query $\varphi$ is defined as a natural language description that corresponds to a subset of tuples $\mathcal{D}_{\varphi}\subseteq\mathcal{D}$, where $\mathcal{D}_{\varphi}=\{t^{\prime}_{1},t^{\prime}_{2},\ldots,t^{\prime}_{k}\}$ represents the target subset intended for retrieval.
-3. The GenAI-powered Monte-Carlo Method
-Nearest-neighbor search in traditional settings is a well-studied line of research where many algorithms and indices such as Voronoi diagrams (Kolahdouzan and Shahabi, 2004), tree data structures (Yianilos, 1993), and local sensitive hashing (Jafari et al., 2021; Huang et al., 2015) have been proposed for efficient exact and approximate query answering.
-In such settings, the data points are specified as points in a $d$ dimensional space, i.e., $t_{i}\in\mathbb{R}^{d}$, where the goal is to find the $k$ nearest neighbors to a query point $q\in\mathbb{R}^{d}$ based on a distance function $dist:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$.
-That is, $k\mbox{-min}_{t_{i}\in\mathcal{D}}dist(q,t_{i})$.
-Such approaches, however, are not directly applicable in our setting because of the following challenges.
-  1. (1)
-    Multi-modal tuples such as images are not represented as points in $\mathbb{R}^{d}$.
-  2. (2)
-    The query is in the form of a (possibly complex) natural language phrase.
-Our key idea for resolving these challenges is to leverage recent advancements in Generative AI (GenAI) and develop a randomized Monte Carlo algorithm that enables applying traditional nearest-neighbor techniques in our setting.
-3.1. Vector Representations
-Addressing the first challenge is possible using the vector representation (aka embedding) of the tuples (Oquab et al., 2014; Kiela and Bottou, 2014).
-A vector representation $\mathcal{E}:dom(t)\rightarrow\mathbb{R}^{d_{\mathcal{E}}}$ is a transformation of the tuples to a high-dimensional vector of numeric values where the semantic similarity of tuples is proportional to the cosine similarity of their vector representations.
-Specifically, given a vector representation $\mathcal{E}^{\ell}$, we present $\mathcal{E}^{\ell}(t_{i})$ as $\vec{v}^{\ell}_{i}=\langle v_{1},v_{2},\cdots,v_{d_{\mathcal{E}^{\ell}}}\rangle$.
+The input query $\varphi$ is defined as a natural language description that corresponds to a subset of tuples $\mathcal{D}_{\varphi}\subseteq\mathcal{D}$, where $\mathcal{D}_{\varphi}=\{t^{\prime}_{1},\cdots,t^{\prime}_{k}\}$ represents the target subset for retrieval.
+[…]
+Vector Representations: A vector representation $\mathcal{E}:dom(t)\rightarrow\mathbb{R}^{d_{\mathcal{E}}}$ is a transformation of the tuples to a high-dimensional vector of numeric values where the semantic similarity of tuples is proportional to the cosine similarity of their vector representations.
+Specifically, given a vector representation (aka embedder) $\mathcal{E}^{\ell}$, we present $\mathcal{E}^{\ell}(t_{i})$ as $\vec{v}^{\ell}_{i}=\langle v_{1},v_{2},\cdots,v_{d_{\mathcal{E}^{\ell}}}\rangle$.
 The distance between two vector embeddings $\vec{v}^{\ell}_{i}$ and $\vec{v}^{\ell}_{j}$ is computed based on their cosine similarity as
-$\delta^{\ell}_{i,j}=\delta(\vec{v}^{\ell}_{i},\vec{v}^{\ell}_{j})=1-\cos\left(\angle(\vec{v}^{\ell}_{i},\vec{v}^{\ell}_{j})\right)$
+$\delta^{\ell}_{i,j}=\delta(\vec{v}^{\ell}_{i},\vec{v}^{\ell}_{j})=1-\cos\left(\angle(\vec{v}^{\ell}_{i},\vec{v}^{\ell}_{j})\right)$ (1)
 Let $\Delta(t_{i},t_{j})$ be the (unknown) semantic distance of the tuples $t_{i}$ and $t_{j}$.
-A vector representation $\mathcal{E}^{o}$ is optimal if it captures the semantic similarity of tuples, i.e., $\delta^{o}_{i,j}\sim\Delta(t_{i},t_{j}),~{}\forall t_{i},t_{j}$. Formally,
-(1) $\Delta(t_{i},t_{j})\geq\Delta(t_{i},t_{k})~{}\Leftrightarrow~{}\delta^{o}_{i,j}\geq\delta^{o}_{i,k}$
+A vector representation $\mathcal{E}^{o}$ is optimal if it captures the semantic similarity of tuples, i.e., $\delta^{o}_{i,j}\sim\Delta(t_{i},t_{j}),~{}\forall t_{i},t_{j}$.
+Formally,
+$\Delta(t_{i},t_{j})\geq\Delta(t_{i},t_{k})~{}\Leftrightarrow~{}\delta^{o}_{i,j}\geq\delta^{o}_{i,k}$ (2)
 In practice, however, vector representations are learned using a universe of available multi-modal data.
@@
 That is, the expected distance between the embeddings of two tuples is proportional to their semantic distance, i.e., $\mathbb{E}\left[\delta^{\ell}_{i,j}\right]\sim\Delta(t_{i},t_{j}),~{}\forall t_{i},t_{j}$.
-3.2. A Baseline Approach
-The state-of-the-art resolution for the second challenge is to develop vector representations that are jointly trained on the multi-modal data and textual data.
-Specifically, CLIP (Radford et al., 2021) is an embedding trained jointly on (images, text) pairs.
-As a result, CLIP transforms text and image to the vector-representation same space.
-Subsequently, given a natural language query, one can compute its CLIP embedding and find its nearest neighbor images in the embedding space.
+A Baseline Approach: A vector representation that is jointly trained on the multi-modal data and textual data transform text and image to the vector-representation space.
+Therefore, given a natural language query, one can compute its embedding and find its nearest neighbor images in the embedding space.
 This approach, however, suffers from two major issues:
-First, the paired text with the images are usually simple sentences (e.g., [A photo of a dog]), missing to fully describe the rich information encoded in the image.
+First, the paired texts with the images are usually simple sentences (e.g., [A photo of a dog]), unable to fully describe the rich information encoded in the image.
 Conversely, our goal is to answer complex natural language queries, describing a scene with multiple objects and their relationships (e.g., [A dog sitting by the fireplace, staring at the cat on the couch]).
-As a result, as we shall verify it in our experiments (§ 6), while such embeddings may perform reasonably on simple natural language queries their performance drops significantly on complex queries.
-Second, as explained in Section 3.1, learned embeddings only preserve semantic similarities in expectation.
+As a result, as we shall verify in our experiments (§ 5), while such embeddings may perform reasonably on simple natural language queries, their performance drops significantly on complex queries.
+Second, as explained earlier, learned embeddings only preserve semantic similarities in expectation.
 As a result, relying solely on one vector representation for query answering may be misleading and inaccurate.
-Therefore, we instead develop a randomized Monte Carlo algorithm that utilized a set of vector representation transformations to develop more accurate query answering.
-3.3. Query Transformation using GenAI
-In traditional nearest neighbor problems both the dataset tuples and the query are represented as points in the same space.
-However, in our problem the query is form of a natural language phrase, while the dataset contains multi-modal tuples.
+Hence, we instead utilize an ensemble (a set) of vector representation transformations, possibly trained on different training datasets, to develop our Monte Carlo algorithm for more accurate query answering.
+3.1 Query Transformation using GenAI
+In traditional nearest-neighbor problems, both the dataset tuples and the query are represented as points in the same space.
+However, in our problem, the query is in the form of a natural language phrase, while the dataset contains multi-modal tuples.
 Let $\varphi$ be the given natural language query, and let $\{\vec{v}^{o}_{1},\cdots,\vec{v}^{o}_{n}\}$ be the vector representations for the tuples in $\mathcal{D}$.
-If we could specify the corresponding vector $\vec{q}^{o}(\varphi)$ for $\varphi$ in the $\mathcal{E}^{o}$ space, we could find the $k$-nearest neighbors of $\varphi$ as
-(2) $k\mbox{-NN}(\varphi,\mathcal{D})=\underset{t_{i}\in\mathcal{D}}{k\mbox{-argmin}}~{}\delta\left(\vec{q}^{o}(\varphi),\vec{v}^{o}_{i}\right)$
+If we could specify the corresponding vector representation $\vec{q}^{o}(\varphi)$ for $\varphi$ in the $\mathcal{E}^{o}$ space, we could find the $k$-nearest neighbors of $\varphi$ as
+$\hskip 170.71652ptk\mbox{-NN}(\varphi,\mathcal{D})=\underset{t_{i}\in\mathcal{D}}{k\mbox{-argmin}}~{}\delta\left(\vec{q}^{o}(\varphi),\vec{v}^{o}_{i}\right)$ (3)
 As a result, if we could generate the synthetic multi-modal tuple $\mathbf{g}_{\varphi}$ for which $\mathcal{E}^{o}(\mathbf{g}_{\varphi})=\vec{q}^{o}(\varphi)$, we could use Equation 3 for query answering.
-Fortunately, recent advancements in generative AI, particularly the emergence of the foundation models (e.g., DALL·E 3) based on guided-diffusion techniques (Zhang et al., 2023; Rombach et al., 2022; Ramesh et al., 2022), have enabled the generation of synthetic multi-modal tuples for complex prompts $p$ (natural language descriptions of the tuple to be generated).
+Fortunately, recent advancements in generative AI, particularly the emergence of the foundation models (e.g., DALLE 3) based on guided-diffusion techniques [80, 57, 54], have enabled the generation of synthetic multi-modal tuples for complex prompts $p$ (natural language descriptions of the tuple to be generated).
 However, due to many reasons such as the randomized nature of the foundation models, the generated tuples [^4] Throughout this paper, we refer to these synthetic tuples as “guide tuples” that represent the natural language query in search space.
@@
 may not be the optimal synthetic tuple $\mathbf{g}_{\varphi}$, but we can assume that the generated tuple is (semantically) in the vicinity of $\mathbf{g}_{\varphi}$.
-Formally, we assume every tuple generated by the foundation model for the prompt $p=toPrompt(\varphi)$ lies in an $\varepsilon$ distance from $\mathbf{g}_{\varphi}$.
-Where $\mathbb{E}[\varepsilon]=0$.
-In other words for a sample tuple $\bar{\mathbf{g}}_{\varphi}$ we have,
-$\mathbb{E}\left[\Delta(\mathbf{g}_{\varphi},\bar{\mathbf{g}}_{\varphi})\right]=0$ $\Rightarrow\mathbb{E}\left[\delta\left(\mathcal{E}^{o}(\mathbf{g}_{\varphi}),\mathcal{E}^{o}(\bar{\mathbf{g}}_{\varphi})\right)\right]=0$
-$\Rightarrow\mathbb{E}\left[\mathcal{E}^{o}(\bar{\mathbf{g}}_{\varphi})\right]=\mathcal{E}^{o}(\mathbf{g}_{\varphi})$
+Formally, we assume every tuple generated by the foundation model for the prompt $p=toPrompt(\varphi)$ lies randomly within a ($d_{\xi^{o}}$-dimensional) ball centered at $\mathbf{g}_{\varphi}$ with the radius $\varepsilon$.
+Therefore, for a sample tuple $\bar{\mathbf{g}}_{\varphi}$ we have,
+$\mathbb{E}\left[\mathcal{E}^{o}(\bar{\mathbf{g}}_{\varphi})\right]=\mathcal{E}^{o}(\mathbf{g}_{\varphi})$
 The final result follows from the principle that if the expected value of the cosine similarity between two vectors is 1, then the vectors are aligned in expectation.
-3.4. The Monte Carlo Algorithm
-We are now ready to develop our algorithm: a Monte Carlo randomized algorithm based on repeated sampling.
+3.2 The Randomized Algorithm
+We are now ready to develop our Monte Carlo randomized algorithm based on repeated sampling.
 Given a natural language query $\varphi$, and at least one foundation model $\mathcal{F}$, at the first step, we generate $m$ synthetic guide tuples $\{\bar{\mathbf{g}}_{1},\cdots,\bar{\mathbf{g}}_{m}\}$.
-As mentioned before, we consider each guide tuple $\bar{\mathbf{g}}_{i}$ as an unbiased sample taken from a distribution with mean $\mathbf{g}_{\varphi}$.
-Instead of relying on the distance estimations one vector representation, we view each embedding as a sample from the space possible transformation functions.
-Hence, we use multiple embedders to minimize the semantic distance estimation error.
-Let $\{\mathcal{E}^{1},\cdots,\mathcal{E}^{l}\}$ be the set of learned vector representations, considered for transformation.
+As mentioned before, we consider each guide tuple $\bar{\mathbf{g}}_{i}$ as an iid (independent and identically distributed) sample from a distribution with mean $\mathbf{g}_{\varphi}$.
+Furthermore, as mentioned earlier, instead of relying on the distance estimations based on one vector representation, we view each embedder $\mathcal{E}^{\ell}$ as a sample from the space possible embedders.
+Hence, we use a set $\{\mathcal{E}^{1},\cdots,\mathcal{E}^{l}\}$ of embedders to minimize the semantic distance estimation error.
 For each embedder $\mathcal{E}^{\ell}$, we assume $\mathbb{E}\left[\delta^{\ell}_{i,j}\right]=\delta^{o}_{i,j},~{}\forall t_{i},t_{j}$.
 Using the generated tuples $\{\bar{\mathbf{g}}_{1},\cdots,\bar{\mathbf{g}}_{m}\}$ and the embedders $\{\mathcal{E}^{1},\cdots,\mathcal{E}^{l}\}$, we estimate the distance of each tuple $t_{i}$ to the query $\varphi$ as
-(3) $\bar{\delta}_{\varphi,i}=\frac{1}{m\,l}\sum_{j=1}^{m}\sum_{\ell=1}^{l}~{}\delta\left(\mathcal{E}^{\ell}(\bar{\mathbf{g}}_{j}),\vec{v}^{\ell}_{i}\right)$
-Theorem 3.1.
+$\hskip 56.9055pt\bar{\delta}_{\varphi,i}=\frac{1}{m\,l}\sum_{j=1}^{m}\sum_{\ell=1}^{l}~{}\delta\left(\mathcal{E}^{\ell}(\bar{\mathbf{g}}_{j}),\vec{v}^{\ell}_{i}\right)$ (4)
+We propose following theorem to show that with a high probability the distance estimations by Equation 4 are near-optimal.
+Theorem 1
+ [^5] The proof is provided in Appendix B.
 Let $\delta_{\varphi,i}=\delta\left(\vec{q}(\varphi),\vec{v}^{o}_{i}\right)$, $\forall t_{i}\in\mathcal{D}$. Given a positive value $\gamma$,
+$\Pr\left(\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\geq(1+\gamma)\right)\vee\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\leq(1-\gamma)\right)\right)\leq\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{3}}+\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{2}}.$
+Following Theorem 1, we can use the estimated distances to identify the $k$-nearest neighbors of $\varphi$ as
+$\overline{k\mbox{-NN}}(\varphi,\mathcal{D})=\underset{t_{i}\in\mathcal{D}}{k\mbox{-argmin}}~{}\bar{\delta}_{\varphi,i}$
+3.3 Practical Optimizations
+In addition to the theoretical results, we implement several practical optimizations to enhance our system’s performance and efficiency.
+Due to space constraints, we provide a concise summary here, with full details elaborated in Appendix C.
+First, we consider aggregating the rankings based on embedders’ trust scores, which are learned on the fly.
+Since embedders are trained on different datasets and employ different architectures, they may perform differently in practice.
+Therefore, we assign a performance (trust) weight to each embedder conditioned on each topic.
+We propose an algorithm that dynamically adjusts the embedders’ weights based on their performance across various topics using users’ feedback, thereby optimizing their contribution to the final aggregated ranking (Appendix C.1).
+To improve the robustness, we incorporate an outlier detection mechanism to discard potentially misaligned or corrupted generated images from the guide set (Appendix C.2).
+Furthermore, we enhance inference speed through two parallel mechanisms: a caching system that leverages the previously issued queries and acquired metadata to expedite similar query responses (Appendix C.3), and a query classifier that distinguishes simple from hard queries, allowing direct retrieval for simple cases without requiring guide image generation (Appendix C.4).
+4 System Overview
+In the preceding section, we presented the theoretical foundations of our approach.
+In this section, we detail how these principles form the design of our image content retrieval system.
+Although our GenAI-powered Monte Carlo method is modality-agnostic, our focus here is on its application to image retrieval.
+Figure 2 illustrates the overall architecture of Needle, which is organized into two primary components: Preprocessing and Inference.
+[^6] System development details is provided in Appendix D
+Preprocessing: The preprocessing phase involves a one-time operation to create vector representations of the input dataset using a set of predefined embedders $\{\mathcal{E}^{1},\cdots,\mathcal{E}^{l}\}$.
+Once the raw image data is processed into vectors, it is indexed within the vector store.
+The vector store provides a maximum inner product search interface with low latency, which is crucial since users expect prompt results from the system.
+Inference: At the query time, we need to generate guide tuples representing the query using foundation models.
+Needle can use multiple (black-box) foundation models to generate guide images, with the number of guide images fine-tuned as a hyperparameter ($m$).
+Different foundation models might produce images with varying levels of quality and realism.
+However, for the purposes of Needle, high-quality, realistic images are unnecessary to achieve accurate search results.
+As a result, lower-quality images are sufficient, as long as they closely resemble the input query.
+In our extended experiments in Appendix E.4, we will explore the impact of different configurations in the image generation process, like the size of the generated images, the number of guide images, the count, and the overall quality of the foundation models in the final results.
+After generating the guide images, Needle uses predefined embedders to transform guide tuples into vector representations and conducts a k-NN lookup for each generated image across all embedders.
+The results are then aggregated into a final ranking, where each embedder’s contribution is calculated according to its current weight.
+This weight reflects the system’s trust in the performance of the corresponding embedder.
+While dynamic adjustment of these weights based on user feedback is supported (Appendix C.1), the default configuration assumes such feedback is unavailable and uses fixed predefined weights, which can be derived from external performance indicators like timm leaderboard scores [^7] https://huggingface.co/spaces/timm/leaderboard.
+This weighting mechanism is general and fully configurable by the user.
+Figure 2: Needle Architecture
+5 Experimental Evaluation
+In this section, we evaluate several key aspects of Needle to demonstrate its capabilities and to evaluate its performance.
+We use multiple benchmark datasets and several baselines for this purpose.
+In the following, we first detail our experiments setup (§ 5.1), followed by a proof-of-concept analysis (§ 5.2) that demonstrates Needle’s efficacy in text-to-image retrieval across diverse object detection (§ 5.2.1) and complex natural language query benchmarks (§ 5.2.2).
+Subsequently, we conduct an ablation study in Appendix E.4 to assess the impact of hyperparameter variations on Needle’s performance.
+Finally, in Appendix E.5 we provide an end-to-end case study, involving human participants’ qualitative feedback on Needle’s responsiveness to arbitrary user-generated queries.
+5.1 Experiments Setup
+Baselines: We evaluate our method using several prominent vision-language models as baselines, representing different architectural and training paradigms.
+These include well-established contrastive learning models such as CLIP [53] and ALIGN [25], the unified multimodal approach of FLAVA [60], and the effective decoupled pipeline combining BLIP for captioning with MiniLM for text embedding [35, 72].
+Each of these models offers a unique perspective on learning joint image-text representations, and together they provide a comprehensive set of comparisons for our work.
+Detailed descriptions of these models and their specific configurations used in our experiments are deferred to Appendix E.1.
+Datasets: We evaluate Needle using a comprehensive set of datasets, categorized into those focusing on object detection and those involving complex natural language queries.
+The object detection benchmarks include Caltech256 [21], MS COCO [38], LVIS [22], and BDD100k [79].
+For complex natural language queries, we utilize COLA [56], Winoground [63], NoCaps [2], and SentiCap [44].
+These datasets collectively provide diverse challenges, from identifying specific objects to understanding nuanced language and handling novel concepts or sentiments, essential for benchmarking zero-shot retrieval performance.
+Detailed descriptions of the datasets are provided in Appendix E.2.
+Evaluation Metrics: We employ distinct evaluation metrics tailored to the characteristics of our two benchmark types: object detection and complex natural language queries.
+For object detection benchmarks, we measure performance using Mean Recall at $k$ (R@k), Precision at $k$ (P@k), Mean Average Precision (MAP), and Mean Reciprocal Rank (MRR).
+For complex natural language query experiments, we report both MRR and Pairing Accuracy (PAcc).
+Detailed definitions and formulas for these evaluation metrics are provided in Appendix E.3.
+Embedders: We use all or a subset of embedders listed in Appendix (Table 3) in our experiments.
+Hardware Configuration: All experiments were conducted on two servers, each equipped with 32 GB of memory, a 12-core CPU, and two Tesla T4 GPUs, running Ubuntu 22.04 LTS.
+Foundation Models and the Monetary Cost: For image generation in our experiments, we employed several state-of-the-art foundation models, namely DALL-E [^8] https://openai.com/index/dall-e-3/, ImagenV3-fast [^9] https://replicate.com/google/imagen-3-fast, Flux-Schnell [^10] https://replicate.com/black-forest-labs/flux-schnell, and RealvisXL-v3.0-turbo [^11] https://replicate.com/adirik/realvisxl-v3.0-turbo.
+These models were chosen for their complementary strengths in generating high-quality images.
+The total monetary cost incurred for utilizing these models was \$257.53 (USD).
+5.2 Proof of Concept
+Figure 3: Illustration of Complex Natural Language Queries extracted from NoCaps [2], Guide images, and CLIP vs. Needle results.
+5.2.1 Task 1: Object Detection
+Table 1: Zero-shot retrieval performance on object detection datasets. For each dataset, we report R@10, P@10, MAP, and MRR. Each cell shows two numbers (All / Hard).
+Caltech256 COCO LVIS BDD
+Baseline R@10 P@10 MAP MRR R@10 P@10 MAP MRR R@10 P@10 MAP MRR R@10 P@10 MAP MRR
+CLIP 0.926 0.150 0.926 0.150 0.939 0.181 0.952 0.193 0.934 0.400 0.934 0.400 0.952 0.477 0.988 1.000 0.177 0.093 0.167 0.083 0.168 0.078 0.316 0.224 0.660 0.033 0.660 0.033 0.670 0.005 0.714 0.048
+ALIGN 0.941 0.375 0.941 0.375 0.947 0.398 0.961 0.541 0.944 0.800 0.944 0.800 0.960 0.895 0.981 1.000 0.215 0.145 0.201 0.130 0.207 0.129 0.379 0.306 0.560 0.000 0.560 0.000 0.573 0.003 0.704 0.014
+FLAVA 0.882 0.258 0.882 0.258 0.903 0.306 0.949 0.491 0.924 0.100 0.924 0.100 0.941 0.281 0.963 1.000 0.185 0.109 0.172 0.097 0.180 0.099 0.321 0.241 0.660 0.033 0.660 0.033 0.698 0.036 0.725 0.083
+BLIP + MiniLM 0.826 0.317 0.826 0.317 0.838 0.372 0.880 0.408 0.941 0.700 0.941 0.700 0.951 0.698 0.975 1.000 0.180 0.115 0.177 0.111 0.179 0.107 0.332 0.260 0.600 0.167 0.600 0.167 0.610 0.144 0.725 0.333
+[…]
+We begin our experiments by evaluating the performance of Needle against the baselines on object detection benchmarks, utilizing the Caltech256, COCO, LVIS, and BDD datasets.
+For each object in these benchmarks, we formulate queries to retrieve them.
+For Needle, we employed three image generation engines (Flux-Schnell, RealVisV3, and ImagenV3-fast) to generate three images per engine, resulting in a total of nine guide images per query, all at MEDIUM quality.
+Additionally, we utilized the ensemble of embedders detailed in Table 3.
+Following the methodology outlined in Appendix E.1, we considered queries with a CLIP Average Precision below 0.5 as the “hard set”.
+Table 1 presents the performance of Needle against the baselines across R@10, P@10, MAP, and MRR metrics.
+Notably, Needle demonstrates superior performance compared to all baselines on both the easy and hard sets.
+Specifically, for the hard set, Needle achieves MAP improvements of 73%, 10%, 93%, and 10% over the second-best baseline on Caltech256, COCO, LVIS, and BDD datasets, respectively.
+These results underscore the substantial advantage of using synthetic data for image retrieval over traditional contrastive learning and image-to-text approaches.
+Furthermore, Figure 3 provides a visual illustration of the performance of Needle (v.s.
+CLIP) and the synthetic guide images generated for each query.
+5.2.2 Task 2: Complex Natural Language Queries
+Table 2: Zero-shot retrieval performance on complex natural language queries benchmarks.
+Cola Winoground SentiCap NoCaps
+Baseline PAcc MRR PAcc MRR MRR MRR
+CLIP 0.578 0.246 0.519 0.426 0.464 0.573
+ALIGN 0.591 0.301 0.554 0.501 0.555 0.704
+FLAVA 0.615 0.336 0.574 0.482 0.546 0.658
+BLIP + MiniLM 0.449 0.195 0.485 0.330 0.331 0.398
+[…]
+Experiments involving complex natural language queries evaluate the system’s ability to handle challenging language and visual nuances across different retrieval tasks.
+We define two types of evaluations:
+First, the Pairing Task assesses the system’s capability to correctly match a caption to its corresponding image when presented with two highly similar images and two corresponding captions that differ only in subtle details.
+Given four possible image-caption pairings, random chance achieves 0.25 Pairing Accuracy (PAcc).
+This task is evaluated using the Winoground [63] and COLA [56] datasets (see Figure 6 for examples from COLA illustrating the challenges of this setup).
+Second, the Full-set Retrieval Task measures the system’s performance when retrieving a specific image from the entire dataset using a single, potentially complex, query.
+For this task, we report the Mean Reciprocal Rank (MRR).
+This task is evaluated on the Winoground [63], COLA [56], SentiCap [44], and NoCaps [2] benchmarks, as they feature challenging captions that require nuanced understanding for successful retrieval from a large pool.
+Table 2 presents the Pairing Accuracy and Mean Reciprocal Rank (MRR) metrics for different aforementioned baselines.
+As illustrated in the table, Needle outperforms the baselines in both PAcc and MRR.
+It is important to note that current foundation models still exhibit limitations in generating images that are fully aligned with an input query.
+In particular, our investigation reveals that these models frequently fail to produce images with the correct compositional ordering.
+We anticipate that further advancements in foundation models will further enlarge the performance gap between Needle and the baselines.
+It is also noteworthy that the BLIP+MiniLM baseline performed significantly worse than the other baselines.
+Initially, we anticipated that this baseline would serve as a strong competitor in handling complex natural language queries, given its design to capture compositional and relational attributes between objects in images; however, the empirical results did not corroborate these expectations.
+In contrast, FLAVA demonstrated exceptionally strong performance in these tests.
+6 Conclusion
+In this paper, we took a step towards addressing the challenge of answering complex natural language queries in multi-modal data settings by proposing a Generative-AI Powered Monte Carlo method.
+Our proposed method utilizes foundation models to generate synthetic guide tuples that capture the complexities of the input query; then, we leverage these guide tuples to represent the natural language query in multi-modal data space and apply traditional nearest neighbor methods.
+Following this method, We developed Needle, an open-source database dedicated to delivering robust, efficient, and high-performance image retrieval [^12] Due to the space constraints, the limitations of our system are discussed in Appendix A..
+Our comprehensive experiments demonstrated that our system significantly outperformed the state-of-the-art techniques in image retrieval.
+While our proposed technique is not limited to image datasets, its development for other modalities, such as audio and video, requires foundation models that are currently not publicly available.
+Upon the availability of such models, we aim to extend the scope of Needle to a wider range of modalities in our future work.
+We appreciate and highly value future contributions from the open-source community—whether through feedback, bug reports, feature requests, or direct code contributions—to help us further enhance and refine the system.
+APPENDIX
+Appendix A Discussions and Limitations
+Reliance on Existing Models: Needle relies on Foundation Models for guide tuple generation and on embedders for generating vector representations of images, which may have their own capabilities, limitations, and inherent biases.
+This reliance presents both strengths and weaknesses in Needle’s architecture.
+The strength lies in its ability to improve performance by upgrading the internal Foundation Models and embedders.
+On the other hand, though, the limitations of these models also limit Needle.
+To mitigate this issue, Needle supports multiple Foundation Models and embedders, enabling it to draw from a broader range of knowledge and reduce the impact of individual model limitations.
+Preprocessing Efficiency: During the preprocessing phase, we employ a predefined set of embedders to generate vector representations of images.
+Needle processes the images in parallel while utilizing the available GPU resources to distribute the workload efficiently.
+Moreover, Needle supports multiple operational modes upon installation.
+For instance, the fast mode utilizes only two embedders which offers preprocessing speeds that are comparable to those of state-of-the-art methods.
+Inference Efficiency: Needle’s inference consists of three main sequential operations: (a) generating guide tuples, (b) k-NN lookup, and (c) ranking aggregation.
+Theoretically, for retrieving $k$ results from a dataset of size $n$ with constants $m$ (number of guide images) and $\ell$ (number of embedders), Step (a) is a constant-time operation.
+Step (b)’s time complexity depends on the underlying vector store algorithm, and in the current implementation, we leverage HNSW data structure index [42], which provides k-NN lookup time complexity of $O(k\log(n))$, while Step (c) uses Fagin’s instance-optimal TA algorithm for rank aggregation [17].
+In practice, however, the bottleneck lies in Step (a), generating guide tuples.
+For instance, in all of our experiments with up to $n=120K$ images, steps (b) and (c) took less than a second, dominated by the image generation time.
+Fortunately, the image generation time is independent of the dataset size $n$ or the output size $k$, and only relies on the number of guide images $m$.
+Moreover, the generation process can run in parallel, significantly improving performance.
+Needle’s current implementation supports both on-device and online image generation models.
+For optimal performance, we recommend using fast on-device generation models to eliminate network latency.
+Moreover, as demonstrated in our ablation study, Needle maintains high accuracy even with lower-quality, lower-resolution images.
+Therefore, employing fast, lower-quality foundation models can significantly accelerate the overall process.
+In addition to these, the builtin Query Complexity Classifier (C.4) enables Needle to generate images only when the query deemed complex for preliminary methods, which reduces the average inference time drastically compared to using Needle for all queries.
+Extensions to Other Modalities: The core idea behind Needle is adaptable to other modalities beyond images, including audio and video.
+However, due to the current limitations in publicly available foundation models, our focus in this paper was on the image data.
+As foundation models for these additional modalities continue to improve and become more accessible, we plan to extend Needle to support them.
+This will allow Needle to operate across multiple data types, unlocking further potential for handling complex, multi-modal queries in the near future.
+Appendix B Proof of Theorem 1
+Theorem 1 Let $\delta_{\varphi,i}=\delta\left(\vec{q}(\varphi),\vec{v}^{o}_{i}\right)$, $\forall t_{i}\in\mathcal{D}$. Given a positive value $\gamma$,
 $\Pr\left(\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\geq(1+\gamma)\right)\vee\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\leq(1-\gamma)\right)\right)\leq\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{3}}+\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{2}}$
-Proof.
-Replacing $\bar{\delta}_{\varphi,i}$ with the right-hand side of Equation 3, we get
+Proof: Replacing $\bar{\delta}_{\varphi,i}$ with the right-hand side of Equation 4, we get
 $\Pr\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\geq(1+\gamma)\right)$ $=\Pr\left(\frac{\frac{1}{m\,l}\sum_{j=1}^{m}\sum_{\ell=1}^{l}~{}\delta\left(\mathcal{E}^{\ell}(\bar{\mathbf{g}}_{j}),\vec{v}^{\ell}_{i}\right)}{\delta_{\varphi,i}}\geq(1+\gamma)\right)$
@@
 $\mathbb{E}\left[\sum_{j=1}^{m}\sum_{\ell=1}^{l}~{}\delta\left(\mathcal{E}^{\ell}(\bar{\mathbf{g}}_{j}),\vec{v}^{\ell}_{i}\right)\right]=m\,l\,\delta_{\varphi,i}$
-Next, applying Chernoff bound, we get
+Next, applying Chernoff bound [49], we get
 $\Pr\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\geq(1+\gamma)\right)\leq\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{3}}$
@@
 $\Pr\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\leq(1-\gamma)\right)\leq\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{2}}$
-Therefore, applying the union bound, we get
+Therefore, applying the Union bound [49], we get
 $\Pr\left(\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\geq(1+\gamma)\right)\vee\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\leq(1-\gamma)\right)\right)\leq\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{3}}+\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{2}}$
-Using the estimated distances, we identify the $k$-nearest neighbors of $\varphi$ as
-$\overline{k\mbox{-NN}}(\varphi,\mathcal{D})=\underset{t_{i}\in\mathcal{D}}{k\mbox{-argmin}}~{}\bar{\delta}_{\varphi,i}$
-4. System Development
-Figure 2. Needle Architecture
-Having discussed the theoretical foundation of our system, we can now explain its architecture and different components in this section.
-While our GenAI-powered Monte Carlo method is modality agnostic, we will explain the development of our system for Image Retrieval.
-The overall architecture of Needle is illustrated in Figure 2.
-The architecture can be divided into two main parts: components involved in Preprocessing [§ 4.1, § 4.2, § 4.3] and Inference [§ 4.4, § 4.5, § 4.6, § 4.7].
-4.1. Preprocessing
-The preprocessing phase involves a one-time operation to create vector representations, also known as embeddings, of the input dataset using a set of predefined embedders.
-Since images are typically regarded as being strongly associated with their representing objects(Zhao et al., 2019), embedding models may underperform when applied to images containing numerous small objects (Liu et al., 2021b).
-To address this limitation, we employ a tiling strategy that divides the input image into multiple tiles based on the density of objects in each region (§ 4.2).
-Each of these tiles is then processed by the embedders, and the resulting embeddings are stored on disk.
-This process can be executed in parallel for each embedder, with performance significantly enhanced by using multiple GPUs to handle tile embedding.
-4.2. Smart Tiling
-Figure 3. An Illustration of Smart Tiling on an image with lots of objects
-Algorithm 1 Smart Tiling
-1:image, d, objects_borders, min_size
-2:image tiles
-3:$(h,w)\leftarrow\text{size of }image$
-4:tiles [$\leftarrow$[]← [ ]
-5:function CountObjects(x, y, w, h)
-6:    return number of unique objects in the region $(x,y,w,h)$
-7:end function
-8:
-9:function Split(x, y, w, h)
-10:    num_objects $\leftarrow$ CountObjects(x, y, w, h)
-[…]
-12:         Append((x, y, w, h)) to tiles
-13:    else
-14:         Compute half_w, half_h
-15:         Calculate objs_left, objs_right, objs_top, objs_bottom
-16:         if vertical split is preferred & aspect ratio is reasonable then
-17:             Split(x, y, half_w, h) $\triangleright$ Split vertically
-18:             Split(x + half_w, y, half_w, h)
-19:         else
-20:             Split(x, y, w, half_h) $\triangleright$ Split horizontally
-21:             Split(x, y + half_h, w, half_h)
-22:         end if
-23:    end if
-24:end function
-25:
-26:Split(0, 0, w, h)
-27:return tiles
-To capture all the details within an image, it is important to focus on densely populated regions.
-Therefore, we should split each image into different tiles based on the density.
-In addition, most embedders are trained on square images (He et al., 2016; Simonyan and Zisserman, 2014; Szegedy et al., 2017) and perform optimally when applied to such shapes (Shahnaz and Mollah, 2022; Azulay and Weiss, 2019), which means excessively tall or narrow tiles should be avoided.
-In order to achieve these two goals, inspired by (Finkel and Bentley, 1974), we propose our Smart Tiling procedure.
-We employ an edge detection technique (Canny Edge Detection (Canny, 1986)) to identify object boundaries within the image and recursively split the image to balance the distribution of objects across tiles as effectively as possible.
-The objective of this algorithm is to divide the image into smaller tiles, where each tile contains no more than a specified number of objects $d$, while maintaining a balanced aspect ratio.
-The algorithm recursively splits the image by first checking if a region contains more objects than the allowed threshold $d$.
-If this condition is met, the region is split either vertically or horizontally.
-The algorithm calculates the object distribution for potential splits and selects the one that provides the most balanced distribution of objects across the two resulting tiles.
-To avoid generating disproportionate tiles, the algorithm also checks the aspect ratio and favors splits that maintain near-square dimensions.
-This recursive process continues until each tile satisfies the object limit constraint or it reaches the minimum acceptable width/height, at which point the tiles are finalized.
-Algorithm 1 presents the pseudocode for our Smart Tiling approach.
-4.3. Vector Store
-Once the raw image data is processed into vectors, it is indexed within the vector store.
-The vector store provides a maximum inner product search interface with low latency, which is crucial since users expect prompt results from the system.
-In our implementation, we employ hnswlib, which offers approximate maximum inner product search capabilities.
-This is consistent with the functionality provided by most vector stores.
-4.4. Query Refinement
-Having discussed the components involved in Preprocessing, we now explore the components used during Inference.
-After receiving the query in natural language text, Needle refines the input query into a suitable prompt.
-The first step is to remove punctuation and redundant characters from the query.
-Additionally, Needle supports more advanced query refinement techniques.
-[…]
-In addition, given that current Foundation Models have limitations in fully understanding some types of prompts (Liu et al., 2022; Feng et al., 2022; Lee et al., 2023; Liu et al., 2024), our system can employ various query refinement techniques (Liu et al., 2021a; Liu and Chilton, 2022; Wang et al., 2023; Witteveen and Andrews, 2022) to paraphrase and reformat the input query.
-These transformations help better align the generated query tuples with the original intent.
-These techniques help overcome prompt misinterpretations and improve the relevance of generated results.
-4.5. Foundation Model and Guide Image Generation
-After refining the query, We need to generate guide tuples representing the query using foundation models.
-Needle can use multiple (black-box) foundation models to generate guide images, with the number of guide images fine-tuned as a hyperparameter, denoted as $m$.
-Foundation models are capable of producing images with varying levels of quality and resolution.
-However, for the purposes of Needle, high-quality realistic images are unnecessary to achieve accurate search results.
-Lower-resolution images are sufficient, provided that they closely resemble the input query.
-In our experiments (§ 6.3), we will explore the impact of $m$, the size of the generated images, and the overall quality of the foundation model in the final results.
-4.6. User Interface and Feedback Process
-Figure 4. Needle User Interface
-The current implementation of Needle offers a web-based user interface that allows users to search based on their queries (Figure 4).
-Users can interact with the system in two modes: Feedback mode (on) or No-Feedback mode (off).
-When Feedback mode is enabled, the guide tuples are presented to the user prior to the search process.
-Users can evaluate the quality and relevance of the generated guide tuples with respect to their query, which can assist Needle in retrieving more relevant images.
-In No-Feedback mode, Needle automatically identifies and discards outlier tuples from the generated images without user intervention (§ 5.2).
-4.7. Embedders and Rankings
-After generating the guide images, Needle utilizes predefined embedders to convert the guide tuples into vector representations and performs a k-NN lookup on the vector stores (§ 4.3).
-The score of each image is determined by the maximum score among all of its tiles (with the entire image also treated as a tile).
-Subsequently, the results are aggregated using our aggregation mechanism (§ 5.1) to produce a final ranking, which is then presented to the user.
-5. Practical Optimization
-In this section, we discuss our practical optimizations to improve the efficiency of our system.
-First, we address the issue of aggregating results from multiple embedders into a unified ranking (§ 5.1), followed by a discussion on detecting anomalies in generated synthetic guide tuples (§ 5.2).
-5.1. Rankings Aggregation and Embedder Trust Mechanism
+Appendix C Practical Optimization
+C.1 Rankings Aggregation and Embedder Trust Mechanism
 Needle has been designed with scalability and flexibility in mind, ensuring it is not restricted to any specific embedder.
@@
 $\sum_{i=1}^{l}w_{i}^{t}=1.$
-5.1.1. Aggregation
-Given the ranked lists from all the embedders and the topic $t$ of the query $\varphi$, the objective is to aggregate these lists into a single final ranking $R_{\varphi}$ by incorporating the topic-specific reliability weights $w^{t}_{i}$.
+Aggregation: Given the ranked lists from all the embedders and the topic $t$ of the query $\varphi$, the objective is to aggregate these lists into a single final ranking $R_{\varphi}$ by incorporating the topic-specific reliability weights $w^{t}_{i}$.
 Let $I_{h}\in\mathcal{D}$ be a tuple returned by at least one of the embedders.
@@
 Let $\mathsf{S}:[k]\rightarrow[0,1]$ be a monotonically decreasing function that specifies the importance of each rank position in the top-$k$.
-In particular, while being agnostic to the choice of $S(.)$, following the literature (Cummins, 2012; Arampatzis and Robertson, 2011; Faria et al., 2010), we adopt the following position-importance function in Needle:
+In particular, while being agnostic to the choice of $S(.)$, following the literature [13, 4, 19], we adopt the following position-importance function in Needle:
 $\mathsf{S}(i)=\begin{cases}\frac{1}{i}&~{}i\leq k\\ 0&~{}\text{otherwise}\end{cases}$
 Then, the aggregated score of $I_{h}$ is computed as:
-(4) $score(I_{h})=\sum_{i:\mathcal{E}^{i}\in\mathbf{E}}w^{t}_{i}\cdot\mathsf{S}(rank(I_{h},R^{i}_{j}))$
+$score(I_{h})=\sum_{i:\mathcal{E}^{i}\in\mathbf{E}}w^{t}_{i}\cdot\mathsf{S}(rank(I_{h},R^{i}_{j}))$ (5)
 The top-$k$ tuples based on Equation 5 are returned as the query output, denoted as $R_{\varphi}$.
-5.1.2. Dynamic Reliability-weight Adjustment
-Suppose after showing the top-$k$ results $R_{\varphi}$ to the user, we receive feedback from him/her, which is a subset of results marked as irrelevant to the query $\varphi$  [^3] We acknowledge that such feedback may not commonly be available after the system deployment.
+Dynamic Reliability-weight Adjustment: Suppose after showing the top-$k$ results $R_{\varphi}$ to the user, we receive feedback from him/her, which is a subset of results marked as irrelevant to the query $\varphi$  [^14] We acknowledge that such feedback may not commonly be available after the system deployment.
 However, crowdsourced feedback on a small set of topic-based benchmarks can be employed to calculate these weights before deploying a new embedder to our system..
 Let $\mathcal{T}=\{I^{\prime}_{1},I^{\prime}_{2},\dots,I^{\prime}_{s}\}$ be the set of irrelevant tuples where $\mathcal{T}\subseteq R_{\varphi}$
-Using this feedback, we update the reliability weights of the embedders for each topic using the Multiplicative Weight Update Method (MWUM) (Arora et al., 2012).
+Using this feedback, we update the reliability weights of the embedders for each topic using the Multiplicative Weight Update Method (MWUM) [5].
 Let $t=t(\varphi)$ be the topic of the query $\varphi$.
 Then the partial loss of the embedder $\mathcal{E}^{i}$ for a topic $t$ based on $\varphi$ is computed as,
-(5) $\Delta loss_{\varphi}(t,i)=\sum_{j:\bar{\mathbf{g}}_{j}\in\mathbf{G}}\sum_{I^{\prime}_{h}\in\mathcal{T}}\mathsf{S}(rank(I^{\prime}_{h},R^{i}_{j}))$
+$\Delta loss_{\varphi}(t,i)=\sum_{j:\bar{\mathbf{g}}_{j}\in\mathbf{G}}\sum_{I^{\prime}_{h}\in\mathcal{T}}\mathsf{S}(rank(I^{\prime}_{h},R^{i}_{j}))$ (6)
 We only penalize those embedders that highly ranked the irrelevant results within the query topic $t$.
@@
 where $\eta\in(0,1)$ is the learning rate controlling how much the trust in each embedder changes over time. Finally, the weights are normalized for topic $t$ to ensure $\sum_{i=1}^{l}w_{i}^{t}=1$.
-5.2. Anomaly Detection in Query Images
+C.2 Anomaly Detection in Query Images
 In addition to rankings aggregation, Needle employs an outlier detection mechanism to identify potentially anomalous guide images based on their embeddings from multiple embedders.
@@
 Our goal is to detect if any guide image $\bar{\mathbf{g}}_{j}$ is an anomaly by analyzing the embeddings across all embedders.
-5.2.1. Dimensionality Reduction using UMAP (McInnes et al., 2018)
-Given the high-dimensional nature of the embeddings, we first apply dimensionality reduction to project the embeddings into a lower-dimensional space.
-For each embedder $\mathcal{E}^{i}$, we reduce the dimensionality of the embeddings $\{\vec{v}^{i}_{1},\vec{v}^{i}_{2},\dots,\vec{v}^{i}_{m}\}$ to a lower dimension $d$ using UMAP (Uniform Manifold Approximation and Projection) (McInnes et al., 2018).
+Dimensionality Reduction using UMAP [45]: Given the high-dimensional nature of the embeddings, we first apply dimensionality reduction to project the embeddings into a lower-dimensional space.
+For each embedder $\mathcal{E}^{i}$, we reduce the dimensionality of the embeddings $\{\vec{v}^{i}_{1},\vec{v}^{i}_{2},\dots,\vec{v}^{i}_{m}\}$ to a lower dimension $d$ using UMAP (Uniform Manifold Approximation and Projection) [45].
 UMAP constructs a weighted $k$-nearest neighbors graph of the high-dimensional points, optimizes the layout in a low-dimensional space, and minimizes the cross-entropy between the two spaces.
@@
 where $d\ll\text{dim}(\vec{v}^{i}_{j})$.
-5.2.2. Outlier Detection using LOF (Breunig et al., 2000)
-After dimensionality reduction, we apply the Local Outlier Factor (LOF)(Breunig et al., 2000) algorithm to identify outliers.
+Outlier Detection using LOF [8] : After dimensionality reduction, we apply the Local Outlier Factor (LOF)[8] algorithm to identify outliers.
 LOF assigns an outlier score to each data point based on the density of its neighborhood.
@@
 This score indicates how much of an outlier $\bar{\mathbf{g}}_{j}$ is in the context of the embeddings from embedder $\mathcal{E}^{i}$, with higher scores representing a greater likelihood of being an outlier.
-5.2.3. Aggregation of Outlier Scores
-The final outlier score of an guide image $\bar{\mathbf{g}}_{j}$ is computed by aggregating the individual scores, weighted with the reliability weight of each embedder:
+Aggregation of Outlier Scores: The final outlier score of an guide image $\bar{\mathbf{g}}_{j}$ is computed by aggregating the individual scores, weighted with the reliability weight of each embedder:
 $S(\bar{\mathbf{g}}_{j})=\sum_{i=1}^{l}w_{i}\cdot\text{LOF}^{i}(\bar{\mathbf{g}}_{j})$
-5.2.4. Anomaly Detection Threshold
-In the end, if the outlier score $S(\bar{\mathbf{g}}_{j})$ of an image $\bar{\mathbf{g}}_{j}$ exceeds a given threshold $\tau$, the image is flagged as an anomaly:
+Anomaly Detection Threshold: In the end, if the outlier score $S(\bar{\mathbf{g}}_{j})$ of an image $\bar{\mathbf{g}}_{j}$ exceeds a given threshold $\tau$, the image is flagged as an anomaly:
 $S(\bar{\mathbf{g}}_{j})>\tau\quad\Rightarrow\quad\bar{\mathbf{g}}_{j}\text{ isan anomaly}$
@@
 In contrast, lower values of $\tau$ force Needle to adhere more closely to the query, leading to higher precision in the results but potentially missing some relevant images.
-6. Experiments
-This section presents results of our experiments designed to evaluate the effectiveness of our proposed system, Needle.
-We begin by outlining the experimental setup (§ 6.1) and proceed with a proof of concept (PoC) (§ 6.2), where we demonstrate Needle ’s ability to retrieve relevant images across different scenarios.
-In the first scenario, we focus on object detection tasks, which may involve identifying either rare (single occurrence) or common objects within the dataset (§ 6.2.1).
-In the second scenario, the goal is to locate a specific image using its natural language description.
-We evaluate how accurately Needle can retrieve the target tuple based on the provided description.
-Following the PoC, we explore the impact of various hyper-parameters on Needle’s performance (§ 6.3) and how can Needle perform well, even with a few low-resolution guide tuples.
-We then analyze the performance of the anomaly detection component embedded within Needle and will compare it with human evaluation results(§ 6.4).
-Finally, we present a case study in which human evaluators compare Needle ’s results with CLIP (§ 6.5).
-6.1. Experimental Setup
-6.1.1. Datasets
-We have used Caltech256 (Griffin et al., 2022), LVIS (Gupta et al., 2019), and the Berkeley DeepDrive (BDD) (Yu et al., 2020) datasets for object detection.
-Caltech256 is a relatively simple benchmark that includes 256 categories of objects with 30,000 images.
-LVIS is a long-tailed benchmark with over 1,300 categories and 100,000 images.
-Due to the fact that some categories in LVIS have fewer than 10 images across the entire dataset, this benchmark is considered particularly challenging for object detection tasks.
-BDD is a large-scale, diverse dataset for autonomous driving, which contains over 100,000 video frames and images annotated for object detection, lane detection, and drivable area segmentation.
-For the targeted queries experiments, we use two image captioning benchmarks: Nocaps (Agrawal et al., 2019) and SentiCap (Mathews et al., 2016).
-Nocaps is a large-scale image captioning dataset that extends the Open Images dataset (Kuznetsova et al., 2020), featuring captions for novel object categories and their relations to each other.
-SentiCap (Mathews et al., 2016), on the other hand, focuses on sentiment-driven image captions, containing images annotated with captions that convey either positive or negative sentiments, making it suitable for evaluating models on emotional and subjective image descriptions.
-6.1.2. Evaluation Metric
-In our experiments, we report Average Precision (AP) as the primary evaluation metric.
-AP is a widely used accuracy metric in information retrieval, as it places greater emphasis on retrieving relevant results earlier in the ranking compared to simpler metrics such as precision or F1-score.
-AP values range from 0 to 1, with an AP of 1 indicating perfect accuracy, meaning all relevant results appear before any irrelevant ones.
-For a given query, AP is calculated as the average of the individual precision scores at each possible recall threshold: $AP=(\sum_{i=1}^{R}P_{i})/R$, where $R$ is the total number of relevant results in the dataset, and $P_{i}$ represents the precision when considering the top $i$ relevant results.
-In addition to reporting the mean Average Precision (mAP) as the final performance indicator for each baseline, we also report Hit Rate for targeted queries.
-Since these queries are designed to find exactly one target image, Hit Rate measures whether the relevant image appears in the results, regardless of its rank.
-The mean Hit Rate (mHR) across all queries is also presented for each system.
-6.1.3. Baselines
-In our experiments, we use the latest and highest-performing version of CLIP, specifically $ViT$-$B/32$, as the baseline model (Radford et al., 2021).
-CLIP (Contrastive Language–Image Pretraining) is a powerful vision-language model developed by OpenAI, designed to learn visual concepts from natural language descriptions.
-It achieves this by jointly training an image encoder and a text encoder using a contrastive learning objective, aligning images and their corresponding textual descriptions in a shared embedding space.
-CLIP has demonstrated strong zero-shot performance across a wide range of visual tasks, making it a robust baseline for image retrieval tasks.
-In all experiments, we will report AP for all queries in the dataset and their “hard subset”. For each dataset, we define the hard subset as the set of tuples for which CLIP AP is less than 0.5.
-6.1.4. Environmental Setup
-All experiments were conducted on a local server equipped with 128 GB of memory, 32 CPU cores, and two NVIDIA GeForce RTX 2080 Ti GPUs.
-Four different Foundation Models were utilized throughout the experiments.
-runwayml-v1.5 and sdxl-turbo-v1.0 were hosted locally on the aforementioned server, while DALL·E 3 and Flux-Schnell were accessed via paid APIs.
-6.2. Proof of Concepts
-6.2.1. Object Detection Experiments
-Table 1. Performance (mAP) Comparison Between CLIP and Needle on Object Detection Benchmark Datasets
-Dataset (#hard/#all) All Queries Hard Subset
-CLIP Needle CLIP Needle
-BDD (2/11) 0.76 0.83 0.01 0.24
-Caltech256 (10/256) 0.94 0.95 0.26 0.59
-LVIS (1049/1203) 0.16 0.31 0.08 0.23
-Average 0.62 0.70 0.12 0.35
-((a)) LVIS
+C.3 Guide Image Caching and Metadata Generation
+((a)) Caltech256
+((b)) LVIS
+((c)) BDD
+Figure 4: Correlation between CLIP cosine similarity and Mean Average Precision (MAP) across various benchmarks.
+The Pearson correlation coefficient ($r$) and its associated p-value are displayed within each subplot.
+The generation of guide images, while instrumental for Needle’s operation, can introduce computational overhead and latency.
+This is particularly true when utilizing large-scale or proprietary generative models that may incur substantial processing times and network data transmission costs.
+To mitigate these inefficiencies and enhance overall system performance and cost-effectiveness, Needle incorporates a novel caching mechanism designed to minimize the frequency of redundant image generation.
+A fundamental challenge in open-domain text-to-image retrieval is the scarcity or inadequacy of rich, comprehensive textual metadata explicitly describing image content.
+As demonstrated in our experiments (Table 1 and Table 2), even advanced image tagging methods often fail to capture the full semantic depth and nuance of visual information.
+Ideally, the availability of precise metadata (e.g., detailed tags or captions) would enable highly efficient text-based similarity searches to retrieve relevant images directly.
+However, the manual generation of such high-quality metadata for large image datasets is generally infeasible.
+We propose an implicit method for metadata generation and caching by leveraging the search queries themselves.
+The underlying premise is that for a given input query $\varphi$, any image $I_{j}$ successfully retrieved by Needle for $\varphi$ is highly likely to be semantically related to $\varphi$.
+This allows us to effectively ’tag’ or associate these retrieved images with the input query $\varphi$, thereby enriching our internal knowledge base.
+This strategy reduces the need for repeated generative processes for similar or semantically related queries, as these tagged images can then be retrieved via a similarity search solely in the text embedding space.
+A critical aspect of this query-based metadata generation is quantifying the confidence in the newly established association between a retrieved image and its tagging query, especially given that the granularity of relevant objects or concepts might differ across various queries.
+To address this, we investigate whether the raw cosine similarity scores produced by a multi-modal embedder, such as CLIP, can serve as a reliable indicator of retrieval confidence.
+We hypothesize that higher similarity scores for relevant items are predictive of a more confident and accurate retrieval for a given query.
+To formally validate this hypothesis, we assess the statistical correlation between an embedder’s mean cosine similarity for relevant images and its retrieval performance, as measured by Average Precision (AP).
+Consider a set of queries $\mathcal{Q}=\{\varphi_{k}\}_{k=1}^{N}$.
+For each query $\varphi_{k}$, let $\mathcal{R}_{\varphi_{k}}$ denote the set of ground-truth relevant images.
+The similarity between vector embeddings of an image $I$ and a query $\varphi_{k}$ using embedder $\ell$ is computed using cosine similarity, denoted as $\cos(\angle(\vec{v}_{I}^{\ell},\vec{v}_{\varphi_{k}}^{\ell}))$.
+We define the mean cosine similarity (MCS) for query $q_{k}$ from an embedder $\ell$ as:
+$\text{MCS}_{\ell}(\varphi_{k})=\frac{1}{|\mathcal{R}_{\varphi_{k}}|}\sum_{I\in\mathcal{R}_{\varphi_{k}}}\cos(\angle(\vec{v}_{I}^{\ell},\vec{v}_{\varphi_{k}}^{\ell}))$ (7)
+We then compute the Pearson correlation coefficient, $r$, between the vector of $\{\text{MCS}_{\ell}(\varphi_{k})\}_{k=1}^{N}$ values and the vector of $\{\text{AP}(\varphi_{k})\}_{k=1}^{N}$ across all queries in $\mathcal{Q}$:
+$r=\frac{\sum_{k=1}^{N}(\text{MCS}_{\ell}(\varphi_{k})-\overline{\text{MCS}})(\text{AP}(\varphi_{k})-\overline{\text{AP}})}{\sqrt{\sum_{k=1}^{N}(\text{MCS}_{\ell}(\varphi_{k})-\overline{\text{MCS}})^{2}\sum_{k=1}^{N}(\text{AP}(\varphi_{k})-\overline{\text{AP}})^{2}}}$ (8)
+A positive Pearson $r$ value close to +1 indicates a strong linear relationship, suggesting that queries yielding higher mean cosine similarity with their relevant images tend to correspond to higher Average Precision.
+This correlation validates the utility of embedder similarity scores as a proxy for retrieval confidence.
+Conversely, a non-significant p-value ($p>0.05$) would imply no statistically reliable linear relationship in the observed data.
+We empirically validate this correlation by calculating $r$ and its corresponding $p$-value for the Caltech256 [21], LVIS [22] and BDD100K [79] datasets using CLIP as the embedder.
+As illustrated in Figure 4, these experiments consistently demonstrate a strong direct correlation between the absolute value of CLIP’s cosine similarity and Mean Average Precision (MAP).
+This observation forms the basis of our image tagging strategy: formally, given an input query $\varphi$ and a fine-tuned threshold $\tau$, Needle finds relevant images, and in a post-retrieval stage, assigns query $\varphi$ to all images $I\in\mathcal{R}_{\varphi}$ for which $\cos(\angle(\vec{v}_{I}^{\ell},\vec{v}_{\varphi}^{\ell}))\geq\tau$.
+We will elaborate on how these generated tags can improve the inference speed of Needle in Appendix C.4.
+C.4 Query Complexity Classifier
+While Needle shows outstanding performance and accuracy in retrieval by using guide image generation in its core methodology, we observe that for many simpler queries, a large computational overhead of image generation may not be necessary to achieve satisfactory results.
+In such cases, computationally efficient retrieval methods, such as conventional contrastive learning approaches like CLIP [53] and ALIGN [25], might provide sufficient performance on their own.
+To optimize performance and resource utilization, we propose a query complexity classifier designed to dynamically assess the complexity of an input query $\varphi$.
+This module is designed to predict the expected Average Precision (AP) of a query’s retrieval if performed solely using these preliminary methods.
+These methods, denoted as =$\mathcal{M}=\{M_{p}\}_{p=1}^{P}$, include established contrastive learning models and our proposed tag-based search mechanism (as introduced in Section C.3), all of which operate without requiring on-the-fly image generation.
+For an input query $\varphi$, each method $M_{p}\in\mathcal{M}$ produces a ranked list of images and their associated cosine similarity scores, denoted as $S_{M_{p}}(I,\varphi)=\cos(\angle(\vec{v}_{I},\vec{v}_{\varphi}))$.
+Specifically, for method $M_{p}$, the retrieval result can be represented as $R_{p}(\varphi)=\{(I_{p,j},S_{M_{p}}(I_{p,j},\varphi))\}_{j=1}^{N}$.
+The design of this complexity classifier prioritizes both computational efficiency and contextual relevance.
+Predicting query complexity solely based on linguistic features of the query text itself can be insufficient, as a query’s actual retrieval effectiveness often depends on the specific content and nuances of the underlying image database being searched.
+Furthermore, for an on-the-fly decision module that must operate upon query arrival, the prediction process should be extremely fast to avoid becoming a bottleneck.
+To this end, we use a feature-based approach where features are derived directly from the preliminary retrieval results, which are already computed, and are designed to capture not only the query’s inherent characteristics but also its interaction with the database.
+This allows the classifier to operate as a lightweight model, analyzing readily available metrics rather than performing complex, potentially time-consuming, end-to-end linguistic analysis of the query text.
+From the top-$K$ retrieved images and their cosine similarity scores for each preliminary method, we extract a comprehensive set of features that collectively indicate the query’s complexity.
+These features form a vector $\vec{f}(\varphi)\in\mathbb{R}^{D}$ and include:
+  • Mean Top-$K$ Cosine Similarity Scores: For each method $M_{p}$, we compute the average cosine similarity score among its top-$K$ retrieved images.
+Let $R_{p,K}(\varphi)=\{I_{p,j}\}_{j=1}^{K}$ denote the set of top-$K$ images retrieved by method $M_{p}$.
+The mean top-$K$ cosine similarity score is:
+    $\bar{S}_{M_{p},K}(\varphi)=\frac{1}{K}\sum_{I\in R_{p,K}(\varphi)}S_{M_{p}}(I,\varphi)$
+    As illustrated in Figure 4, a higher $\bar{S}_{M_{p},K}(\varphi)$ generally suggests a higher expected AP for that query.
+  • Top-$K$ Inter-Method Overlap Coefficients: To quantify the consensus among different preliminary methods, we compute overlap metrics between their top-$K$ retrieved sets.
+For any pair of distinct methods $(M_{p},M_{q})$ from $\mathcal{M}$, the Jaccard index is used:
+    $J(R_{p,K}(\varphi),R_{q,K}(\varphi))=\frac{|R_{p,K}(\varphi)\cap R_{q,K}(\varphi)|}{|R_{p,K}(\varphi)\cup R_{q,K}(\varphi)|}$
+    High overlap indicates strong agreement across methods, typically characteristic of queries that yield higher AP values with preliminary models.
+  • Confidence Deviation: We also use measures of consistency in cosine similarity scores within the top-$K$ results for each method.
+The standard deviation of the top-$K$ cosine similarity scores, $\sigma_{M_{p},K}(\varphi)$, serves as an indicator of confidence deviation:
+    $\sigma_{M_{p},K}(\varphi)=\sqrt{\frac{1}{K-1}\sum_{I\in R_{p,K}(\varphi)}(S_{M_{p}}(I,\varphi)-\bar{S}_{M_{p},K}(\varphi))^{2}}$
+    A lower standard deviation suggests higher confidence and uniformity in cosine similarity scores for a given query, which is often observed for queries achieving higher AP.
+This constructed feature vector $\vec{f}(\varphi)$ is then used as an input into a pre-trained regression model $\mathcal{C}:\mathbb{R}^{D}\to[0,1]$.
+For the purpose if real-time operation upon query arrival, $\mathcal{C}$ is designed as a very simple and lightweight machine learning model (e.g., a linear regression model, or a shallow neural network) that operates efficiently on the derived feature vector $\vec{f}(\varphi)$, rather than performing computationally expensive end-to-end text analysis.
+The regression model outputs a predicted Average Precision score, $\text{AP}_{\text{pred}}(\varphi)$, for the given query using preliminary methods.
+The model $\mathcal{C}$ is trained using supervised learning on a dataset of queries, where the ground-truth target for each query is its actual Average Precision (AP) score when retrieved using the preliminary methods.
+A predefined threshold $\gamma_{AP}\in[0,1]$ is applied to the predicted Average Precision score $\text{AP}_{\text{pred}}(\varphi)$ to define a dynamic routing strategy:
+  • If $\text{AP}_{\text{pred}}(\varphi)\geq\gamma_{AP}$, the query is categorized as "simple." In this case, Needle implements a short-circuiting mechanism, directly returning the early retrieval results obtained from the preliminary ensemble.
+This bypasses the computationally intensive image generation process, significantly reducing latency and computational cost.
+  • If $\text{AP}_{\text{pred}}(\varphi)<\gamma_{AP}$, the query is categorized as "complex." In this scenario, Needle proceeds with its full pipeline, including guide image generation and subsequent Monte Carlo sampling, to perform a more complex retrieval.
+Appendix D System Development
+D.1 Development Objectives
+As one of our primary objectives, we aim to develop Needle as an open-source deployment-ready application that both researchers and developers can easily adopt.
+Specifically, we considered a set of design objectives, outlined in the following; later, in § D.2, we describe how our system implementation fulfills these objectives.
+Database Functionality and CLI Integration: Needle should be easily installable across various operating systems and offer an intuitive command-line interface (CLI) for efficient user interaction.
+Deployment Readiness and Versatility: Needle should be engineered for immediate deployment, operating out-of-the-box on a wide range of system configurations, including resource-limited personal computers.
+Configurability and Abstract Database Modes: Users should have the freedom to select from multiple operational modes (fast, balanced, accurate) that balance retrieval accuracy and latency according to their specific application needs.
+Additionally, Needle should allow for customization beyond these predefined modes.
+GPU Acceleration and Scalability: Given the computational intensity of image indexing, Needle should automatically leverage available GPU resources and support workload distribution across multiple GPUs to ensure scalable performance.
+Modular Embedder Architecture: As embedders are central to Needle’s functionality, the architecture should facilitate easy integration and upgrading of new embedder models, ensuring the system remains at the forefront of technological advancements.
+Multi-Directory Management: Needle should be capable of managing multiple image directories concurrently.
+It should allow users to add, remove, or toggle directories for search while indexing new directories in the background without interrupting ongoing queries.
+Robustness and Consistency Maintenance: Needle should perform consistency checks upon restart after downtime to synchronize any database changes.
+It should continuously monitor designated directories for modifications, ensuring the database index remains aligned with the file system.
+Integration with External Image Generators: To facilitate the generation of guide images, Needle should provide connection wrappers for leading proprietary and open-source image generation services (e.g., DALL-E, Google Imagen, Replicate).
+Furthermore, it should allow users to integrate any image generator that complies with the specified request schema.
+Flexible Output Formats and Enhanced Usability: Recognizing that Needle may serve as a component within larger pipelines, it should support multiple output formats, including JSON, YAML, and human-readable text.
+Additionally, built-in query previews should be provided to improve the overall user experience.
+((a)) needlectl Commands Overview and Adding a Directory
+((b)) Image Generator/Image Directory Configurations
+((c)) Running queries using needlectl and Needle Results
+((d)) Needle Query Results Builtin Preview Page for [a train] in BDD dataset [79] – 102 out of 70K images include a train
+Figure 5: Screenshots of key interfaces and features of Needle.
+D.2 System Details
+Having defined our development objectives, we can now delve into the Needle’s implementation details.
+Architecture: Needle by nature requires multiple services to work together.
+It needs a service for handling Image Generation, a service for handling Directory Integration and Indexing Progress, and a service for storing calculated embeddings and kNN lookup.
+In order to handle these multiple sources of responsibilities, we have designed Needle with a microservice architecture, delegating correlated responsibilities to their respective services.
+Needle’s services are:
+Backend: This microservice serves as the gateway to the core functionalities of Needle.
+It encapsulates the primary operations of the system, exposing APIs such as /search, /directory, and /query.
+Additionally, it manages the initialization of embedders and oversees the image indexing process.
+Database: The backend container needs persistent storage to store directory structures, indexing progress, and other critical metadata.
+Needle is agnostic to the choice of the database, but in the current implementation, we use PostgreSQL [^15] https://www.postgresql.org/ for this goal.
+Vector Store: Needle requires a vector store that supports embedding, indexing, and inner product searches.
+This service is used to store the outputs of various embedders and to efficiently execute k-nearest neighbor (kNN) searches for a given query.
+In the current implementation, we have used Milvus [^16] https://milvus.io/ for this purpose.
+Image Generator Hub: This service manages the image generation process by providing wrappers for renowned services such as DALL-E, Imagen, and Replicate.
+It also supports custom image generation models that conform to its defined schema.
+Moreover, the hub offers fallback mechanisms and priority settings to maximize robustness and ensure reliable image generation.
+Command-Line Interface: Needle requires an intuitive interface to work with backend service and apply user’s requests. This service is the exposed UI of Needle to the end user.
+Since the individual services have their own dependencies and requirements, we containerized all core services using docker to mitigate dependency conflicts and make deployment easier across various operating systems.
+We employ docker compose to orchestrate these containers, establish a shared network, and configure the individual services.
+Additionally, to eliminate the need for direct Docker manipulation, we developed a Command-Line Interface (CLI), needlectl, as a standalone binary.
+needlectl installs into the user’s path and provides seamless access to the Needle core.
+Technical Stack and Frameworks: Backend and ImageGeneratorHub services built using FastAPI framework [^17] https://fastapi.tiangolo.com/.
+For GPU-intensive tasks, PyTorch [^18] https://pytorch.org/ is leveraged to distribute workloads across available GPUs, ensuring efficient processing.
+Embedder management is streamlined via integrations with Hugging Face and the timm library [^19] https://github.com/huggingface/pytorch-image-models, facilitating the initiation and upgrading of embedding models.
+The CLI client (needlectl) is developed using the Typer [^20] https://typer.tiangolo.com/ library to build intuitive command-line applications, while textual [^21] https://textual.textualize.io/ is utilized to manage terminal user interfaces (TUIs).
+Standalone Python libraries are packaged with PyInstaller [^22] https://pyinstaller.org/en/stable/, simplifying the distribution and execution.
+Installation, upgrading, and uninstallation processes are automated using Bash scripts, ensuring a smooth user experience.
+Containerization and orchestration are achieved with Docker [^23] https://www.docker.com/ and Docker Compose, respectively.
+For documentation, we employ mdBook [^24] https://rust-lang.github.io/mdBook/ to create comprehensive and well-structured guides.
+Finally, all of the build, development, and versioning processes are automated via CI/CD pipelines using GitHub Actions [^25] https://github.com/features/actions.
+Command-Line Interface and APIs: Needle offers a robust CLI that enables seamless interaction with its core services.
+The needlectl is provided as a standalone binary, allowing users to execute commands directly without requiring additional dependency installations.
+This design ensures that non-expert users can easily operate Needle.
+The CLI commands and options have been carefully designed to be intuitive and straightforward.
+The command structure adheres to the following format:
+needlectl [--global options] [component] [action] [--action options]
+This structure supports clear and consistent command usage with the following key components (some illustrative examples are provided in Listing LABEL:listing:1):
+  • service: for managing Needle-related tasks
+  • directory: for handling image directory operations
+  • query: for managing active queries
+  • generator: for interfacing with image-generation functionalities
+needlectl service start     # Starts the Needle services
+needlectl service log       # Retrieves service logs
+needlectl directory add /path/to/image/dir --show-progress
+# Adds an image directory for indexing and monitoring, displaying progress
+needlectl --output json query run "a wolf" --num-engines 2 --num-images-to-generate 4 --image-quality LOW
+# Executes a query using two generators, each generating four low-quality images, with JSON output
+Installation and Deployment:
+Needle is installed using a single one-liner command that executes a Bash script.
+This script automatically verifies system prerequisites and checks for GPU availability.
+It detects the current operating system and downloads the appropriate configuration.
+During installation, users are prompted to select a preferred database mode—Fast, Balanced, or Accurate—which optimizes various parameters such as the number of embedders, default image resolution for generation, default number of generators, number of images generated per query, and HNSW[42] index construction parameters (e.g., M and ef).
+Finally, the installation script guides users through the initial steps to start and utilize the Needle service.
+Table 3: Embedder Models and Their Weights
+Architecture Model Name Weight
+eva[18] eva02_large_patch14_448 0.8497
+regnet[75] regnety_1280 0.8235
+dinoV2[51] vit_large_patch14_reg4_dinov2 0.8235
+CLIP[53] vit_large_patch14_clip_336 0.8146
+convnextV2[73] convnextv2_large 0.8184
+bevit[7] beitv2_large_patch16_224 0.7660
+Appendix E Additional Experiments
+E.1 Details of Baseline Models
+1.
+CLIP [53]: Developed by OpenAI [^26] https://huggingface.co/openai/clip-vit-base-patch32, this model uses a ViT-B/32 image encoder and a Transformer-based text encoder.
+It learns a shared embedding space via contrastive learning on large-scale image–text pairs, and its widespread use and benchmarking make it a standard reference.
+We use CLIP as our standard baseline and designate categories where CLIP’s average precision (AP) is below 0.5 as "hard categories"—instances where CLIP struggles to produce satisfactory results.
+This filtered subset is then used to demonstrate the enhanced capability of alternative baselines in effectively handling rare or difficult queries.
+2.
+ALIGN [25]: Originally developed by Google, ALIGN (A Large-scale ImaGe and Noisy-text Embedding) employs a dual-encoder architecture similar to CLIP, using separate encoders for images and text trained with a contrastive loss.
+However, while CLIP (in our clip-vit-base-patch32 variant) uses a Vision Transformer for image encoding, ALIGN typically uses a CNN image encoder along with a Transformer for text.
+Moreover, the original ALIGN was trained on a much larger, noisier dataset (over 1.8 billion image–text pairs) than CLIP.
+In our experiments, we use an open-source version [^27] https://huggingface.co/kakaobrain/align-base that preserves the core architecture but is trained on different data, achieving even better accuracy on some benchmarks.
+3.
+FLAVA [60]: [^28] https://huggingface.co/facebook/flava-full a unified multimodal model developed by Facebook that jointly learns representations for images and text.
+Unlike CLIP, which primarily relies on a contrastive approach applied to image-text pairs, and ALIGN, which emphasizes scaling up representation learning using noisy text supervision with separate encoders, FLAVA adopts a holistic pre-training strategy.
+It integrates both unimodal and multimodal objectives, enabling FLAVA to capture richer semantic interactions across modalities.
+4.
+BLIP + MiniLM: This pipeline approach first converts images into descriptive captions using BLIP [35] [^29] https://huggingface.co/Salesforce/blip-image-captioning-base—a state-of-the-art image captioning model from Salesforce known for its high-quality, informative captions.
+The generated captions are then transformed into embeddings with MiniLM [72] [^30] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2, a robust text encoder widely adopted in industrial applications.
+This decoupled strategy not only leverages mature text retrieval systems but also enables independent optimization of the captioning and text-embedding stages.
+As a result, it has become a strong baseline and is employed in many commercial products, offering enhanced interpretability and scalability compared to end-to-end models like CLIP.
+E.2 Dataset Details
+((a)) Blue napkin to the right of white plate
+((b)) White napkin to the left of blue plate
+Figure 6: Illustration of COLA compositional benchmark queries, using each image description, where the goal is to identify the correct image
+Object Detection Datasets:
+  1. 1.
+    Caltech256 [21]: Contains 30,607 images spanning 256 object categories. Its diverse object classes make it a useful benchmark for assessing the robustness of detection and retrieval models.
+  2. 2.
+    MS COCO [38]: With over 118K images and 80 object categories, MS COCO is widely used for object detection and segmentation.
+Its comprehensive annotations and varied scene compositions provide a challenging testbed.
+  3. 3.
+    LVIS [22]: offers instance segmentation with a long-tail distribution of over 1,200 categories.
+Its focus on rare and fine-grained objects is critical for evaluating retrieval performance on less frequent classes.
+  4. 4.
+    BDD100k [79]: This dataset is tailored to urban driving scenarios with detailed annotations, making it valuable for testing retrieval in real-world, dynamic contexts.
+Natural Language Query Datasets:
+For evaluating the performance of various baselines on complex natural language queries, we focus on scenarios where models must differentiate between two highly similar images that exhibit nuanced differences.
+In these settings, the model is expected to identify the correct image by carefully attending to subtle details in both the visual content and the corresponding query.
+To rigorously assess these capabilities, we utilize several compositional benchmarks:
+  1. 5.
+    COLA [56]: As a benchmark for compositional text-to-image retrieval, this dataset features nuanced captions that require the system to distinguish between multiple similar images by capturing subtle semantic and spatial details.
+  2. 6.
+    Winoground [63]: Designed to evaluate visio-linguistic alignment, Winoground presents pairs of images and captions that differ only in their compositional structure.
+This dataset challenges models to accurately map nuanced language to the corresponding image, serving as a stringent test of fine-grained retrieval performance.
+  3. 7.
+    NoCaps [2]: A large-scale benchmark for novel object captioning, featuring images from Open Images with human-generated captions.
+It is crucial for evaluating a model’s ability to handle zero-shot retrieval on images containing objects not commonly found in standard captioning datasets.
+  4. 8.
+    SentiCap [44]: Derived from datasets like MS COCO, SentiCap provides images with captions specifically annotated for positive or negative sentiment.
+It allows for evaluating retrieval systems’ ability to understand and retrieve images based on subjective or emotional language cues, moving beyond purely factual descriptions.
+E.3 Evaluation Metrics
+We use different evaluation metrics for object detection and complex natural language queries benchmarks.
+In object detection benchmarks, we define each benchmark dataset $\mathcal{D}$ as a set of images $\mathcal{D}=\{t_{1},t_{2},\ldots,t_{p}\}$, where each image $t_{i}$ comprises a set of objects $t_{i}=\{o_{1},o_{2},\ldots,o_{s}\}$, with $o_{j}$ representing an individual object within image $t_{i}$.
+For each benchmark, we compute the union of all objects present across all images, denoted as $\bigcup_{i=1}^{n}t_{i}=\bigcup_{i=1}^{n}\{o_{1},o_{2},\ldots,o_{s}\}$, and utilize each unique object $o_{j}$ within this union as a query $q$ for each retrieval engine.
+For a given query $q$, we retrieve a ranked list of files =$\mathcal{F}=\{f_{1},f_{2},\ldots,f_{n}\}$ ($n=60$), representing the retrieved images.
+We construct a list of relevance scores =$\mathcal{R}=\{r_{1},r_{2},\ldots,r_{n}\}$, where $r_{i}=1$ if $q$ is present in image $f_{i}$, and $r_{i}=0$ otherwise.
+We start list indexing from 1, so by definition $\mathcal{R}[j]=r_{j}$.
+We assume access to a function $C(.)$, which gets an object as input and returns the number of images that this object appears in.
+We are interested in retrieving at most 10 positive instances.
+Therefore, we define the effective number of positives as: $ep=\min(10,C(q))$.
+Given $\mathcal{R}$ and $ep$, We assess the performance of the retrieval using the following metrics:
+Mean Recall at $k$ (R@k): $r@k=\frac{\sum_{i=1}^{k}\mathcal{R}[i]}{ep}$.
+A higher r@k indicates that a larger proportion of the relevant images (up to 10) are retrieved in the top $k$ positions, we report mean r@k (R@k) which is the average of r@k over all objects in the dataset.
+Precision at $k$ (P@k): $p@k=\frac{\sum_{i=1}^{k}\mathcal{R}[i]}{k}$.
+A higher p@k reflects that a greater proportion of the retrieved images are relevant, which is crucial when the user examines only the top results, same as before, we report the averaged out P@k over all object in the dataset.
+Mean Average Precision (MAP): measures the average precision (AP) over all queries.
+$AP=\frac{1}{ep}\sum_{i=1}^{n}\left(\frac{\sum_{j=1}^{i}\mathcal{R}[j]}{i}\right)\cdot r_{i},$
+where $r_{i}=\mathcal{R}[i]$ and the effective number of positives is defined as $ep=\min(10,C(q))$.
+The overall mean average precision (MAP) is then the average of $AP$ over all objects in the dataset.
+MAP provides a single-figure measure that considers both the precision and recall across the entire ranked list, rewarding systems that rank relevant images higher.
+Mean Reciprocal Rank (MRR): computes the reciprocal rank for each query and then averages these values over all queries.
+For a given query $q$, let $k_{q}$ be the smallest index such that $\mathcal{R}[k_{q}]=1$, indicating the rank of the first relevant image.
+The reciprocal rank for $q$ is defined as:
+$RR(q)=\begin{cases}\frac{1}{k_{q}},&\text{if }\exists\,k\text{ with }\mathcal{R}[k]=1,\\[5.0pt] 0,&\text{otherwise.}\end{cases}$
+The Mean Reciprocal Rank (MRR) is the average value over all queries. MRR emphasizes the importance of retrieving at least one relevant image as early as possible for ranking.
+For the complex natural language query experiments, we evaluate our method using both MRR and Pairing Accuracy.
+In these experiments, the baseline is given two images, $t_{1}$ and $t_{2}$, along with two captions, $c_{1}$ and $c_{2}$, and each caption must be assigned to its corresponding image correctly.
+Pairing Accuracy is defined as the number of correct assignments divided by the total number of queries.
+Note that the pairing accuracy for a random baseline is 0.25% (chance of choosing the correct combination over all possible combinations).
+E.4 Hyper-parameter Ablation Study
+((a)) BDD
 ((b)) Caltech256
-((c)) BDD
-Figure 5. Distribution of Average Precision for different queries for CLIP and Needle, split into hard and easy subsets.
-In these experiments, we demonstrate the effectiveness of Needle in basic object detection tasks.
-We utilize state-of-the-art benchmark datasets, LVIS (Gupta et al., 2019), Caltech256 (Griffin et al., 2022), and BDD (Yu et al., 2020), for this evaluation.
-For each query, the number of retrieved images is set to 60 ($k=60$), and we aim to identify 10 relevant tuples within the results ($R=10$).
-Two local Stable Diffusion models are employed as Foundation Models: runwayml-v1.5 and sdxl-turbo-v1.0 with guide image resolution of $768\times 768$.
-For each model, two guide images are generated, resulting in a total of four guide images per query ($m=4$).
-Table 1 presents a comparative analysis of performance between CLIP and Needle across the different datasets.
-Notably, Needle consistently demonstrates superior performance across all categories, in particular, for the hard subsets, Needle achieves around 200% improvement in AP compared to CLIP on average, highlighting its effectiveness in addressing more challenging object detection scenarios.
-Figure 5 shows the AP distribution of CLIP and Needle for different datasets.
-Needle demonstrates significant improvements in AP, even for queries where CLIP initially performs very poorly, with AP close to zero.
-It is important to note that all generated guide images are produced using local Stable Diffusion models, which generate lower quality images than commercial alternatives.
-This result highlights that Needle imposes much looser constraints on the quality of guide tuples than humans do.
-Even when the guide images are distorted or lack realism, Needle can robustly retrieve relevant images using these guide tuples.
-6.2.2. Targeted Queries
-Figure 6. Illustration of different specific queries, the generated guide images and the comparison of CLIP and Needle results
-Table 2. Performance (mAP, mHR) Comparison Between CLIP and Needle on Targeted Queries Benchmark Datasets
-Dataset (#queries) mAP mHR
-CLIP Needle CLIP Needle
-Nocaps (100) 0.11 0.37 0.55 0.73
-SentiCap (100) 0.24 0.34 0.71 0.77
-Average 0.18 0.36 0.63 0.75
-Needle excels at handling targeted queries and identifying rare tuples.
-We demonstrate this by running two image captioning benchmarks, Nocaps (Agrawal et al., 2019), and SentiCap (Mathews et al., 2016), using both CLIP and Needle.
-For each experiment, the systems are provided with a query that describes a specific image in the dataset, with the search engines returning 10 images.
-For this task, we enhance Needle with two more powerful and precise Foundation Models: DALL·E 3  [^4] https://openai.com/index/dall-e-3/ and Flux-Schnell  [^5] https://replicate.com/black-forest-labs/flux-schnell.
-These models are selected due to the need for guide images that accurately represent the detailed input queries, as these queries include significant complexity and detail.
-In addition to reporting mean Average Precision (mAP), we also evaluate the mean Hit Rate (mHR).
-Hit Rate is particularly useful in cases where the number of retrieved images is small, and the focus is on finding the correct image rather than its specific ranking.
-As shown in Table 2, Needle significantly outperforms CLIP in locating the target image.
-Needle successfully retrieves the correct image in over 75% of queries within the top 10 results, marking around 20% improvement over CLIP’s performance in mHR and 100% improvement in mAP.
-Figure 6 presents four different queries from the Nocaps (Agrawal et al., 2019) dataset, along with the generated guide images for Needle and the top-1 results from both CLIP and Needle.
-As illustrated, CLIP frequently overlooks key aspects of the input prompt and focuses on a single keyword rather than capturing the full intent of the query.
-In contrast, Needle effectively captures all inherent complexity of query within its guide tuples and identifies the target image using only two guide images.
-6.3. Parameter Tuning
-((a)) Runwayml (Caltech256)
-((b)) sdxl-turbo (Caltech256)
-((c)) Runwayml (LVIS)
-((d)) sdxl-turbo (LVIS)
-Figure 7.
-Heatmaps of Needle mAP for two different datasets (Caltech256 and LVIS) using two local Foundation Models (Runwayml and sdxl-turbo) while altering number of Generated Images $m$ and Generated Image Size
-In this section, we investigate the effect of varying the number of generated guide images ($m=\{1,2,4\}$), the image resolution ($\{512\times 512,768\times 768,1024\times 1024\}$), and the performance of different foundation models (runwayml, sdxl-turbo) on Needle’s effectiveness.
-The experiments are conducted on both the Caltech256 and LVIS datasets.
-Figure 7 illustrates the mAP values for different combinations of these parameters.
-As a general trend, increasing the number of images and the resolution improves mAP performance.
-However, there are deeper insights to consider: first, sdxl-turbo exhibits more consistent and slightly better performance across different image sizes compared to runwayml; second, while increasing the number of generated images $m$ improves results, the improvement diminishes and stabilizes after a certain point.
-A similar trend is observed for image sizes.
-These results suggest that Needle can maintain high performance even with a limited number of generated images and lower resolution inputs.
-6.4. Feedback Mode
-((a)) LVIS
-((b)) BDD
-Figure 8. Comparison of AP distribution for Needle on Feedback mode on and off
-CLIPNeedleBothNeither002020202040404040606060608080808010010010010023.2323.2323.2323.2352.5252.5252.5252.5214.1514.1514.1514.1510.110.110.110.1Percentage (%)
-Figure 9. Case study results
-Table 3. Comparison of Needle’s performance (mAP) for different datasets with Feedback mode enabled (ON) and built-in Anomaly detection (OFF)
-Dataset (#hard/#all) All Queries Hard Subset
-OFF ON OFF ON
-BDD (2/11) 0.82 0.84 0.22 0.25
-LVIS (84/100) 0.28 0.30 0.19 0.21
-Average 0.55 0.57 0.21 0.23
-In this subsection, we investigate the impact of enabling the feedback procedure in Needle on mAP.
-A human evaluator was employed to verify the quality and alignment of the generated images with the input query prior to retrieving results.
-For evaluation, we used LVIS and BDD datasets, generating four tuples per query ($m=4$) with an image size of $768\times 768$, utilizing the runwayml foundation model.
-We compare the mAP of Needle’s built-in anomaly detection method (§ 5.2) to the performance when human feedback is incorporated.
-Figure 8 presents the AP distribution for both baselines.
-As shown, there is a slight improvement when feedback mode is enabled, primarily raising queries with $AP\approx 0.4$ in LVIS and $AP\approx 0.8$ in BDD toward higher APs.
-This indicates that: (a) Needle’s built-in anomaly detection method performs effectively in practice, achieving results comparable to a human evaluator, and (b) human intervention is most beneficial for queries that already have a moderate AP, whereas queries for which foundation models fail to generate suitable images do not significantly improve with feedback alone.
-6.5. Case Study
+((c)) COCO
+((d)) LVIS
+Figure 7: Illustration of how varying number of guide images and number of embedders involved affects Needle MAP
+In this section, we examine the impact of each hyper-parameter on Needle’s performance.
+Generated Guide Images and Embedders Count: We investigate the impact of the number of images generated per engine ($m$) and the number of embedders used ($\ell$) on the performance of Needle.
+We vary $m$ over $\{1,2,3\}$ and $\ell$ over $\{1,2,4,6\}$, and conduct experiments with these configurations on the BDD, Caltech256, COCO, and LVIS datasets.
+While incrementing the value of $\ell$, we add embedders sorted by their weight in descending order (Higher performing embedders are being used first).
+Figure 7 shows the mean average precision (MAP) achieved for each combination.
+In general, increasing both $m$ and $\ell$ leads to improved MAP.
+Notably, even the lowest configuration ($m=1$, $\ell=1$) outperforms most of the baselines evaluated in Table 1 across different benchmarks.
+From this ablation study, we conclude that consensus among embedders can lead to higher reliability and improved performance in Needle.
+Note that Needle supports an arbitrary number of embedders, allowing users to increase this number until performance saturates.
+This experiment also demonstrates that increasing the number of generated images enhances performance; specifically, the performance gap between $m=1$ and $m=2$ is larger than that between $m=2$ and $m=3$.
+This indicates that Needle can achieve good results with a small number of images and does not require extensive image generation to capture all relevant complexities.
+Foundation Models: Needle supports image generation using various open-source and proprietary foundation models.
+These models differ in the quality of generated images, realism, and their alignment with the input prompt.
+In this section, we investigate two primary aspects.
+First, we assess the performance of each individual foundation model when used on its own.
+Second, we examine how combining multiple foundation models can improve performance by providing diverse representations of the input query.
+Our experiments involve four foundation models: DALL-E3 (proprietary), Imagen3-fast (proprietary), Flux-Schnell (open-source), and RealViSXL (based on Stable-diffusionXL; open-source).
+We evaluate their performance on object detection tasks using the BDD, Caltech256, and COCO datasets, as well as on complex natural language queries using the Winoground and COLA benchmarks.
+The inclusion of tests on complex natural language queries is motivated by the higher demands for prompt alignment and image quality in such scenarios.
+We calculate MAP for Object detection and Accuracy for Complex natural language benchmarks as the performance metric.
+Figure 8 demonstrates how each foundation model performed in different benchmarks.
+Notably, the two open-source models, Flux-Schnell and RealVisXL, outperformed proprietary models by a huge margin in Complex queries, indicating they have better prompt alignment and understanding compared to DALL-E3 and Imagen3-fast.
+DALL-E3 performed worse than others in most of the benchmarks, and RealVisXL demonstrated to be a reliable option for most cases.
+Figure 9 illustrates that increasing the number of foundation models leads to an improvement in the overall performance of Needle.
+Notably, the addition of further foundation models yields more improvement in performance for complex natural language queries than for object detection benchmarks.
+This outcome is likely attributable to the increased likelihood of generating images that accurately capture the nuanced details of the input prompt, thereby facilitating the correct identification of the target image.
+In contrast, for object detection tasks—where the objects are generally simpler to generate—the primary benefit of including additional foundation models is an increase in recall due to the diverse representations of the target object across various formats and styles.
+0.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111BDDCaltech256COCO0.690.690.690.690.910.910.910.910.960.960.960.960.730.730.730.730.920.920.920.920.950.950.950.950.610.610.610.610.910.910.910.910.930.930.930.930.680.680.680.680.940.940.940.940.960.960.960.96MAPObject Detection BenchmarksImagen3-fastFlux-schnellDALL-ERealVisXL0.40.40.40.40.450.450.450.450.50.50.50.50.550.550.550.550.60.60.60.60.650.650.650.65COLAWinoground0.60.60.60.60.430.430.430.430.620.620.620.620.510.510.510.510.610.610.610.610.440.440.440.440.610.610.610.610.530.530.530.53AccuracyComplex NLQ Benchmarks
+Figure 8: Comparison of Performance of Different Image Generators on Object Detection and Complex Natural Language Query benchmarks.
+11112222333344440.70.70.70.70.80.80.80.80.90.90.90.91111#Foundation ModelsMAPObject Detection BenchmarksBDDCOCOCaltech256
+11112222333344440.50.50.50.50.550.550.550.550.60.60.60.60.650.650.650.650.70.70.70.7#Foundation ModelsAccuracyComplex NLQ BenchmarksWinogroundCOLA
+Figure 9: The effect of number of Foundation Models on the Performance of Needle for different Benchmarks
+SMALLMEDIUMLARGE0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.8Image SizeMAPBDDDALL-EFlux-SchnellRealVisXLSMALLMEDIUMLARGE0.80.80.80.80.90.90.90.91111Image SizeCOCO
+Figure 10: Illustration of variation of Needle MAP based on size of guide image for different foundation models
+Image Quality: Foundation models usually support generating images in different resolutions.
+In this ablation study, we investigate the effect of image resolution (size) of guide images on Needle’s performance.
+We define three supported image sizes for each foundation model, SMALL, MEDIUM and LARGE.
+The exact amount of pixels is determined based on the foundation model design.
+We generate 3 images per query and utilize Flux-Schnell, RealVisXL and DALL-E for this study (ImagenV3-fast does not support different images sizes as input).
+Figure 10 illustrates the MAP performance of various foundation models on the LVIS and COCO benchmarks across different image quality settings.
+It is evident that foundation models are typically fine-tuned for a specific image resolution, which in turn yields the best quality and prompt alignment.
+In our experiments, DALL-E demonstrated a consistent improvement in performance with increasing image sizes, whereas Flux-Schnell and RealVisXL exhibited a less pronounced correlation between performance and image size, particularly between the MEDIUM and LARGE settings.
+CLIPNeedleBothNeither0020202020404040406060606023.2323.2323.2323.2352.5252.5252.5252.5214.1514.1514.1514.1510.110.110.110.1Percentage (%)
+Figure 11: Results from the human evaluation case study on the LVIS dataset.
+Among the queries with a clear preference, Needle was chosen in approximately 70% of cases, underscoring its practical effectiveness in retrieving relevant images.
+Query Complexity Classifier: This section details the empirical investigation into the impact of our proposed Query Complexity Classifier on the overall efficiency and retrieval performance of Needle.
+The primary objective is to demonstrate how dynamically routing queries based on their predicted complexity can significantly reduce computational overhead without compromising retrieval effectiveness.
+((a)) Comparison of Actual AP and Predicted AP using Needle Query Complexity Classifier
+((b)) Comparison of Inference Speed and MAP using different baselines and benchmarks
+Figure 12: Illustration of how Needle Query Complexity Classifier improves Needle inference speed
+For these experiments, the query complexity classifier is implemented as a Linear Regression model, trained to predict the expected Average Precision (AP) of a given query when processed by preliminary retrieval methods.
+We specifically leverage features derived from two highly efficient preliminary methods: CLIP [53] and ALIGN [25].
+The features utilized for this prediction, as detailed in Section C.4, include mean top-$K$ cosine similarity scores, top-$K$ inter-method overlap coefficients, and confidence deviation for both CLIP and ALIGN.
+The model was trained on a combined dataset derived from the training splits of the Caltech256 [21] and LVIS [22] benchmarks.
+The ground-truth target for training was the actual AP achieved by the preliminary methods for each query.
+Figure 12(a) showcases the performance of our trained classifier in predicting AP for Caltech256.
+The primary benefit of the query complexity classifier is its ability to dynamically route queries, thereby enhancing the inference speed of Needle.
+To quantify this improvement, we conducted experiments comparing the average inference speed and overall retrieval accuracy (Mean Average Precision, MAP) of Needle, both with and without the classifier, against several established baselines.
+The baselines include: CLIP, ALIGN, and BLIP+MiniLM.
+Our evaluation was performed across diverse benchmarks, including Caltech256, LVIS, and BDD100k [79].
+Figure 12(b) illustrates a comprehensive comparison of average inference speed and retrieval accuracy for all evaluated baselines and configurations of Needle.
+As one can observe from the results, employing our proposed query complexity classifier enables Needle to achieve significantly faster inference speeds with only a minimal or negligible loss in overall retrieval performance.
+This is primarily attributed to the classifier’s ability to accurately identify "simple" queries, for which the computationally intensive guide image generation process can be entirely bypassed.
+Furthermore, it is important to note that Needle’s efficiency is designed to progressively increase over time through its caching mechanism.
+As textual metadata for different images are extracted from user’s queries (C.3), the percentage of cache hits will naturally rise, further reducing the overhead of on-the-fly image generation and consequently enhancing inference speed.
+E.5 Case Study
 To further validate Needle’s practical applicability, we conducted a case study with human evaluators.
@@
 For each query, we retrieved 20 relevant images from both CLIP and Needle, displaying the results side by side with randomized orientations for each query.
-Evaluators were then asked to select the more relevant results, choosing from four options: ”Left is better,” ”Right is better,” ”Both are good enough,” and ”Neither is good enough.”, then we recorded the engine behind that option as preferred engine.
+Evaluators were then asked to select the more relevant results, choosing from four options: "Left is better," "Right is better," "Both are good enough," and "Neither is good enough".
+Then, we recorded the engine behind that option as the preferred engine.
 A total of 20 human evaluators participated in the study, contributing 100 queries.
-The queries ranged from simple ones like ”Celery stew” to more complex descriptions such as ”A person walking on the sidewalk of a river in Fall at Chicago.”, Figure 9 presents the results of this user study.
+The queries ranged from simple ones like "Celery stew" to more complex descriptions such as "A person walking on the sidewalk of a river in Fall at Chicago.", Figure 11 presents the results of this user study.
 If we consider only the queries where one engine is preferred over the other, Needle is chosen as the better engine in almost 70% of cases.
-7. Related Work
-Our paper relates to work in the following areas:
-Multi-modal Data Retrieval: Multi-modal data retrieval, which integrates various data types (e.g., text, images, video, and audio), has attracted significant attention.
-Early methods focused on shared representations by mapping modalities into a common latent space using canonical correlation analysis (CCA) (Rasiwasia et al., 2010; Wang et al., 2015b) and kernel-based techniques (Jia et al., 2019).
-Recent approaches leverage deep learning models, including cross-modal encoders (Wang et al., 2015a; Liu et al., 2017) and joint embedding networks (Balaneshin-kordan and Kotov, 2018; Rubio et al., 2017).
-Pre-trained transformers like CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) highlight the effectiveness of combining vision and language models for cross-modal tasks.
-Advanced techniques, such as attention mechanisms (Yang et al., 2021; Tang et al., 2023) and contrastive learning (Pham et al., 2024), further enhance retrieval by capturing fine-grained relationships.
-Feedback mechanisms (Moll et al., 2023; Zhou and Huang, 2003) and user preferences (Zhang and Zhang, 2010; Chen et al., 2020) improve multi-modal search results.
-Research also extends to other modalities, such as text-video retrieval (Li et al., 2019; Markatopoulou et al., 2017; Torabi et al., 2016; Mithun et al., 2018; Miech et al., 2019; Luo et al., 2021) and text-audio retrieval (Xin et al., 2023; Chechik et al., 2008; Koepke et al., 2022).
-For a detailed survey of cross-modal retrieval methods and future directions, see (Li et al., 2023).
-GenAI in Data Management: The intersection of Generative AI (GenAI) and data management has emerged as a promising area, driving innovations in how data is generated, managed, and utilized.
-GenAI models, such as GPT (Achiam et al., 2023) and Stable Diffusion (Rombach et al., 2022), have been employed for synthetic data generation to augment datasets (Tanaka and Aranha, 2019; Shahbazi et al., 2024), improve model training (Le et al., 2017; Tremblay et al., 2018), and address data scarcity or imbalance issues (Erfanian et al., 2024).
-In data cleaning and integration, these models assist with tasks like imputation (Villaizán-Vallelado et al., 2024), and anomaly detection by generating plausible data patterns (Ye et al., 2024).
-Recent works also explore the role of GenAI in automating query generation (Trummer, 2022, 2023; Li and Xie, 2024) and data summarization (Dar et al., 2024), enhancing the efficiency of database systems and analytics platforms.
-Natural Language Query Processing: The evolution of natural language to structured query translation has advanced from rule-based systems to sophisticated large language models (LLMs).
-Early methods used template matching and predefined rules (Androutsopoulos et al., 1995; Li and Jagadish, 2014), which were inflexible and required frequent manual updates.
-Machine learning approaches introduced more adaptability by learning query patterns from annotated datasets (Ma et al., 2020; Zhang et al., 2019), though they struggled with complex language constructs.
-Deep learning models, especially transformers like BERT (Devlin, 2018) and GPT (Brown, 2020), enabled semantic understanding and laid the foundation for text-to-SQL systems (Li et al., 2024).
-Modern LLMs, such as PaLM (Chowdhery et al., 2023) and GPT-4 (Achiam et al., 2023), generate context-aware queries across multiple domains with minimal supervision.
-Retrieval-augmented generation (RAG) further enhances query precision by leveraging external knowledge (Lewis et al., 2020).
-Ongoing efforts focus on fine-tuning LLMs to improve accuracy and efficiency in converting natural language to structured queries (Trummer, 2022, 2023; Li and Xie, 2024).
-8. Discussions and Limitations
-8.1. Reliance on Existing Models
-Needle extensively relies on Foundation Models for guide tuple generation and on embedders for generating vector representations of images.
-Both these components have their own capabilities, limitations, and inherent biases.
-This reliance presents both strengths and weaknesses in Needle’s architecture.
-The strength lies in the ability to easily improve performance by upgrading the internal Foundation Models and embedders.
-However, the limitations of these models also become limitations for Needle.
-To mitigate this, we designed Needle to support multiple Foundation Models and embedders, enabling it to draw from a broader range of knowledge and reduce the impact of individual model limitations.
-8.2. Efficiency
-We assess the efficiency of Needle across two phases: Preprocessing and Inference.
-Preprocessing: In the preprocessing phase, two highly parallelizable tasks are performed: tiling and embedding creation.
-Tiling is a fast operation, but it can linearly increase the time required for embedding creation.
-To improve time efficiency, one can increase the minimum tile width and height, thereby reducing the number of tiles.
-The time required for embedding creation can be significantly improved by utilizing multiple GPUs.
-Since the preprocessing phase is a one-time operation that can be easily parallelized and fine-tuned, it does not present a major concern for Needle’s overall performance.
-Inference: Needle’s inference consists of three main sequential operations: (a) generating guide tuples, (b) k-NN lookup, and (c) ranking aggregation.
-For an input dataset of size $n$, retrieving $k(1\leq k\leq n)$ results with constants $m$ (number of guide images) and $l$ (number of embedders), (a) is a constant-time operation with respect to $n$ and $k$.
-Step (b)’s time complexity depends on the underlying vector store algorithm, and for HNSW data structure (Malkov and Yashunin, 2018), is $O(k\log(n))$.
-Step (c) has a time complexity of $O(k)$ using Fagin’s Algorithm (Fagin et al., 2001).
-In our experiments, where the number of tuples (tiles) is at most 400,000 for any k, (b) k-NN lookup and (c) the rankings aggregation take less than one second using current approximate vector stores.
-The bottleneck of the entire process lies in (a), generating guide tuples.
-The good news is that this operation is independent of the dataset size $n$ and number of images to retrieve $k$, and only relies on the number of guide images $m$.
-Moreover, the generation process can run in parallel, significantly improving performance.
-To summarize, for small values of $n$, Needle’s inference time may be slower than other methods due to the constant time required for guide image generation.
-However, as $n$ increases, the effect of step (a) diminishes, and Needle’s efficiency becomes comparable to other methods.
-8.3. Extensions to Other Modalities
-The core idea behind Needle is adaptable to various modalities beyond images, including audio, video, and others.
-However, due to the current limitations in publicly available foundation models for audio and video generation, our focus in this paper has been restricted to the image modality.
-As foundation models for these additional modalities continue to improve and become more accessible, we plan to extend Needle to support them.
-This will allow Needle to operate across multiple data types, unlocking further potential for handling complex, multi-modal queries in the near future.
-9. Conclusion
-In this paper, we took a step towards addressing the challenge of answering complex natural language queries in multi-modal data settings by proposing a Generative-AI Powered Monte Carlo method.
-Our proposed method utilizes foundation models to generate synthetic guide tuples that capture the complexities of the input query; then, we leverage these guide tuples to represent the natural language query in multi-modal data space and apply traditional nearest neighbor methods.
-Following this method, We implemented Needle for image retrieval.
-We proposed different solutions for practical challenges in our system’s development.
-Our comprehensive experiments demonstrated that our system, Needle, significantly outperforms the state-of-the-art techniques in image retrieval.
-While our proposed technique is not limited to image datasets, its development for other modalities, such as audio and video, requires foundation models that are currently not publicly available.
-Upon the availability of such models, we aim to extend the scope of Needle to a wider range of modalities in our future work.